### v2

# ACPO: AI-Enabled Compiler-Driven Program Optimization

Abstract.
The key to performance optimization of a program is to decide correctly when a certain transformation should be applied by a compiler. This is an ideal opportunity to apply machine-learning models to speed up the tuning process; while this realization has been around since the late 90s, only recent advancements in ML enabled a practical application of ML to compilers as an end-to-end framework.
This paper presents ACPO: AI-Enabled Compiler-driven Program Optimization; a novel framework to provide LLVM with simple and comprehensive tools to benefit from employing ML models for different optimization passes. We first showcase the high-level view, class hierarchy, and functionalities of ACPO and subsequently, demonstrate a couple of use cases of ACPO by ML-enabling the Loop Unroll and Function Inlining passes and describe how ACPO can be leveraged to optimize other passes. Experimental results reveal that ACPO model for Loop Unroll is able to gain on average 4% compared to LLVM’s O3 optimization when deployed on Polybench. Furthermore, by adding the Inliner model as well, ACPO is able to provide up to 4.5% and 2.4% on Polybench and Cbench compared with LLVM’s O3 optimization, respectively.
30(2,0.5)
LLVM Dev 2023 Presentation: https://youtu.be/jzT-tar_X0U
1. Introduction
State-of-the-art compilers are built with sophisticated optimization pipelines that result in robust, reproducible (Ivie and Thain, 2018), and reliable code generation (Aho et al., 1986; Hall et al., 2009). The LLVM compiler framework is no exception with its multi-layer design, where passes in the mid-end can perform analyses or transformations on a code region, one after another (Lattner, 2008). LLVM passes operate on code regions with varying levels of granularity, including loops, functions, call graphs, and modules. Compiler engineers have continuously improved the quality of the optimization passes in LLVM, and have derived a set of fixed-length optimization pipelines, known as standard optimization levels, i.e., O{1,2,3,s,z}, which are designed to improve the code size or performance (to varying degree) of the programs to which they are applied. At the highest optimization level available, i.e. O3, more than 60 unique transformation passes exist in LLVM’s pipeline (Ashouri et al., 2017).
Leveraging machine learning (ML) for compilers has become an extensively researched topic in the past years. In many instances, the authors showed experimental results that outperformed LLVM’s O3 (Agakov et al., 2006; Chen et al., 2012; Park et al., 2013; Ashouri et al., 2016, 2017; Cummins et al., 2017; Georgiou et al., 2018; Haj-Ali et al., 2020; Ashouri et al., 2022) or Deep Neural Networks (DNN) kernel schedulers (Patabandi et al., 2021; Patabandi and Hall, 2023) and lately, using Large Language Models (LLMs) (Cummins et al., 2023; Gong et al., 2024). Machine learning approaches normally rely on data generated by an autotuner (Ansel et al., 2014; Kalyan et al., 2019), which helps with exploring an optimization space formed by the number of parameters in the decision set. Alternatively, by the growth of applications of Reinforcement Learning (RL), as opposed to generating offline training data by an autotuner, authors have explored their action space, online, and during the training process of the RL agent. This is done by means of feeding back a reward value to update the agent’s policy (Trofin et al., 2021; Cummins et al., 2021b; Ashouri et al., 2022; Wang et al., 2022a). To characterize the software being compiled for the ML model, program representation methods are employed, which generally fall into three categories: using static features (Fursin et al., 2011; Ashouri et al., 2022), using dynamic features (Cavazos et al., 2006; Hoste and Eeckhout, 2007), and graph-based characterization (Park et al., 2012; Ben-Nun et al., 2018; Alon et al., 2019; Cummins et al., 2021a). In recent years, there have been considerable advancements in the space, especially in the graph-based program representation techniques (Guo et al., 2020; Cummins et al., 2021a) and token-based identifiers (Wang et al., 2021), which are outside the scope of this work.
The vision for ACPO is to enable the use of Machine Learning in a compiler, in a way that is modular, extendable, and easy to use. ACPO aims to allow for easy collaboration between compiler developers and ML architects by use of clear and simple APIs that enable the collection of relevant data for training purposes, providing inputs to an ML model during inference, and separating the compiler from the model and ML framework to enable model development and testing within the compiler context with minimal intrusion into the compiler code base. This enables the compiler engineers to optimize passes with an easy-to-use set of tools and limited ML knowledge. At the same time, once an ML model is integrated into the compiler, it can be easily optimized and redesigned with no changes to the compiler code base, effectively enabling ML architects to contribute to compiler improvement. This is possible due to ACPO’s separation of the ML framework and the compiler, making it possible for the compiler to work just as easily with any ML model as it is to work with any ML framework, e.g., TensorFlow (Abadi et al., 2016), PyTorch (Paszke et al., 2017), etc, and even other representation mechanism (Fursin et al., 2008; Hoste and Eeckhout, 2007; Cummins et al., 2021a). To this end, ACPO proposes the following contributions:
As a framework, it provides a comprehensive set of tools, libraries, and algorithmic methods by enabling compiler engineers with a straightforward and user-friendly interface to instantiate ACPOModel classes to replace LLVM’s existing hard-coded heuristics.
ML APIs and the compiler are seamlessly connected, but at the same time, they are not interdependent; changing compiler versions, ML models, or ML frameworks won’t break the functionality of ACPO. As long as the inputs/outputs match, users can easily revise the ML side without the need to rebuild LLVM after every change.
Showcasing the benefits of ACPO, we demonstrate two different scenarios:
Loop Unroll Pass — Building both the interface and the ML model. Loop Unroll pass can be optimized from scratch by adding the necessary code to instantiate ACPOLUModel from ACPOModel and also by providing an ML model to outperform LLVM’s unrolling decisions at O3 on Polybench and a number of other benchmarks.
Function Inlining Pass — Building the interface and leveraging an existing ML model. Function Inlining pass can be optimized by leveraging an existing ML model, i.e., MLGOPerf (Ashouri et al., 2022), using ACPO framework. We do that by creating an ACPOModel instance, called ACPOFIModel, and ensuring its inputs and outputs match the existing model so it can be simply invoked at the appropriate point in the compilation process. We demonstrate the ease of integration and versatility of ACPO framework and that how it can be used to connect existing models to the new frameworks. Experimental results on Cbench reveal that the performance of ACPOFIModel is comparable with the reported results in the original work while benefiting from a multi-layer design of ACPO on the separation between the ML APIs and the compiler.
The rest of the paper is organized as follows. Section 2 discusses the state of the art. Section 3 gives details on our proposed framework. Under Sections 4.1 and 4.2, we showcase how Loop Unroll and Function Inlining passes can benefit by instantiating ACPOLUModel. Section 5 showcases the experimental results including the individual and the combined results of deploying both models and Finally, in Section 6 we discuss challenges and propose directions for future work.
2. Related Work
A recent survey on the compiler autotuning and optimization (Ashouri et al., 2018) reveals that there have been a vast number of approaches to employ ML techniques for automatically constructing optimization heuristics. These approaches mainly focused on two major problems of compiler optimization: (1) the optimization selection (Bodin et al., 1998; Cooper et al., 2002; Agakov et al., 2006; Fursin and Temam, 2009; Park et al., 2013; Ashouri et al., 2016) and (2) the phase ordering problem (Kulkarni and Cavazos, 2012; Ashouri et al., 2017; Nobre et al., 2018; Huang et al., 2019; Mammadli et al., 2020; Cummins et al., 2021b; Wang et al., 2022a). Further classification of the two aforementioned problems and by the type of the approach, provides insights for categorizing the works into three groups: (1) autotuning-based approaches (Fursin et al., 2011; Chen et al., 2012; Ashouri et al., 2013; Ansel et al., 2014; Kalyan et al., 2019; Cummins et al., 2021b; Wang et al., 2022b; Park et al., 2022), (2) end-to-end frameworks of optimizations (Fursin et al., 2008; Cummins et al., 2017; Haj-Ali et al., 2020; Cummins et al., 2021b; Rotem and Cummins, 2021; Das et al., 2020; Mannarswamy and Das, 2022; Ashouri et al., 2022; Wang et al., 2022b), and (3) ML-guided approaches (Rotem and Cummins, 2021; Trofin et al., 2021; Ashouri et al., 2022).
Autotuning-based approaches leverage an iterative methodology to wrap around the compiler and to speed up the search in finding better solutions to optimizing a single or multiple passes, often constructing a new sequence of optimization to outperform standard optimization levels by means of phase-ordering or optimization selection (Agakov et al., 2006; Ashouri et al., 2016; Cummins et al., 2021b; Wang et al., 2022a). This can be done in a coarse (Kulkarni and Cavazos, 2012; Ashouri et al., 2016, 2017) or fine-grained manner (Kalyan et al., 2019; Phothilimthana et al., 2021; Cummins et al., 2021b). End-to-end approaches often provide a framework that can be paired or built with the compiler and to offer boosted search or ordering of optimizations (Cummins et al., 2017; Haj-Ali et al., 2020; Mannarswamy and Das, 2022). ML-guided approaches build the decision-making capabilities of ML models into single or multiple passes that can act as a standalone ML-enabled pass. Their corresponding ML models can be either built with the compiler using an Ahead-of-Time (AOT) compilation or be used at runtime using an alternative interprocess communication method, e.g., Python named pipes. All things considered, several recent works can be included in multiple classes, therefore, the lists here are not mutually exclusive.
Cummins et al. (Cummins et al., 2021b) propose CompilerGym, a fine-grained RL environment to speed up the search for tuning a single or multiple compiler passes defined as action space. The authors leverage OpenAI’s Gym (Brockman et al., 2016) library to provide a flexible environment to tackle the selection and the phase-ordering of optimization. The results showcase the benefits of using their RL agent to further reduce the code size of Cbench benchmarks. Contrary to this work, we are proposing an ML-guided approach to provide a framework to optimize passes. Additionally, our objective is to improve the performance as opposed to reducing the code size of a binary.
Trofin et al. (Trofin et al., 2021) propose MLGO that uses ML to advise the Inliner as to whether or not a call site should be inlined, in order to minimize the size of the generated code. MLGO framework compiles the ML model into C++ libraries using Ahead-of-Time (AOT) compilation which becomes part of the compiler at build time. Additionally, the authors have extended the code base to provide optimization for the register allocation problem. Later, MLGOPerf (Ashouri et al., 2022), extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on a number of benchmarks, e.g. SPEC CPU 2006 and Cbench.
In this work, we focus on the hard problem of ML performance optimization, where the key challenge is the development of model architecture and the features used to effectively capture key performance use cases and to facilitate the generation of robust ML models. In the process, we develop a scalable, extendible, end-to-end ML-guided framework to enable easy use of ML within optimization passes in compilers such as LLVM. We demonstrate the effectiveness and the generality of our framework by applying ML-enabling Loop Unroll and Function Inlining passes. We show how ML can easily and effectively be incorporated into a compiler while maintaining separation from the underlying ML frameworks. This framework is easily applicable to other passes and tasks within the compiler domain.
3. Proposed Methodology
Figure 1 presents the high-level design of our infrastructure. There are multiple components involved with the ACPO architecture: Autotuner, compiler, ACPOModel abstraction, low/high-level ML APIs, and ML Framework. For the most part, the autotuner is a wrapper around the compiler and its hooks in the compiler are used to generate model training data. During the inference flow, i.e., when ML models are used for decision making within the compiler, the autotuner is not used and we can think of the compiler as being at the top of the stack.
The ACPO inference flow begins when an ML-enabled pass is invoked and creates an instance of ACPOModel. The instance of ACPOModel aggregates appropriate features from the code region being analyzed and specifies the kind of output required from the ML model. The ACPOModel object then transfers the features, along with information about the model being invoked, to our ML framework via a set of remote procedure call APIs. The APIs provide the ability to send requests, including loading an appropriate model into the ML framework, executing an inference call, and returning the result of the inference back to the compiler. The compiler then applies transformations to the input program as prescribed in the inference output it receives.
In the following sections, we describe in more detail the training and inference flows, and then elaborate on each component of the ACPO framework to show how they fit with the training and inference flows.
Autotuner
The autotuner parameters can be either coarse- or fine-grained. The former includes the selection and the phase-ordering of different optimization passes as a black box, as well as optimization parameters that apply to the whole translation unit, while the latter facilitates the tuning of optimization parameters for individual code regions of a program, e.g. instructions, loops, functions, etc., independently of each other. In this work, we perform fine-grained autotuning with an autotuner (Kalyan et al., 2019), which is built on top of OpenTuner (Ansel et al., 2014). This allows us to generate meaningful training data and target labels that capture the effect of optimization parameters on code region types of interest.
3.1. ACPO Training
A key aspect of creating an ML-driven compiler is the training of models on sample data to ensure that the compiler makes decisions consistently with good results contained in the training data. To do this, we opted for a two-step approach: generate data based on a set of benchmarks the model is expected to reasonably reproduce and then train the model with the collected data. The data generation step is shown in Figure 2a.
To generate data for training purposes we rely on the autotuner. To do this efficiently, we enabled what we refer to as fine-grained autotuning to control each decision of a given transformation pass independently of one another. For example, in the case of loop unrolling, we enabled the autotuner to search for a good unroll factor for each loop, instead of applying an unroll threshold that would be specified on the compiler command line and hence applied globally on all loops. With the help of fine-grained autotuning, we are able to specify which pass we wish to tune and apply this kind of tuning to a set of benchmarks. At the compilation stage, with the instructions from the autotuner, the compiler proceeds to optimize an application and eventually reaches the ML-enabled pass for which the autotuner is currently providing decisions.Note that ACPO supports leveraging different scopes of features as long as it is supported by the compiler and it is up to the developer to decide which ACPO pass might benefit from which scope during the process of design and implementation. The features are passed with std::vector of pairs to the ML side. This process can also be scaled to an arbitrary number of features and feature types. The transformation decision as prescribed by the autotuner is applied to the code region (e.g. the autotuner can set the unroll factor of a loop to a specific value), and then the compilation continues normally until it reaches another code region that requires a decision from the ML model, at which time the process is repeated. Once the compilation is completed, the autotuner runs the program and collects a measurement for performance feedback (the simplest metric being the execution time of the program). We use the O3 compilation as our baseline and measure impact in terms of speedup relative to the O3 compilation.
3.2. ACPO Inference
The compilation process involving ML inference is depicted in Figure 2b. The input program is supplied to Clang to be translated into LLVM Intermediate Representation (IR). The LLVM IR is then passed through a sequence of passes and eventually reaches a pass that uses ACPOModel for its profitability analysis. The pass proceeds normally to confirm the legality of the transformation; when it needs to decide the profitability of the transformation, the ML framework is engaged. The first step in the inference process is the instantiation of the ML model, via the ACPOModel abstraction. This step sets up the interface to the underlying ML framework, such as TensorFlow (Abadi et al., 2016). Upon successful framework initialization, ACPOModel collects features required for inference and sends them to the ML framework. The ML framework is then instructed to run inference on the model, and the inference result is returned to the compiler. If the inference result is such that the transformation should be applied to a given code region, the transform pass then continues and applies the transformation. Otherwise, the pass gracefully rejects the option to transform the code and proceeds to either the next code region to be analyzed or terminates. In some cases, the inference result can contain a numeric value, e.g. the predicted optimal unroll factor for a given loop. To accomplish the tasks of training and inference we used several key components, which we describe in greater detail in the following subsections.
3.3. Compiler
The compiler of choice is LLVM (Lattner, 2008) version 12, which at a high level comprises the C/C++ frontend Clang, the mid-end optimizer, and a target-specific backend to generate executable code. The compiler first translates a program into LLVM IR, and then proceeds to optimize the IR via a sequence of passes. The passes are applied in sequence and at various scopes (module, function, loop, basic block). Once a pass has finished transforming a code region, appropriate internal data structures, including IR, are updated and the pass either proceeds to the next code region or terminates to enable the next pass to proceed. Once the compilation is done, an executable is produced that can be run and measured. We generally refer to the sequence of passes as a pipeline, and we focus on keeping the pipeline unchanged. This is important because empirical evidence shows that the order of passes and composition of the compiler pipeline affect program performance (Ashouri et al., 2017). In this work we focus on improving decision making within a pass using ML, so changes to the pipeline ordering and optimization selection of passes are orthogonal and outside the scope of the work.
3.4. ACPOModel
The ACPOModel class is responsible for storing model-specific information and leveraging ML APIs to performing inference. The header file code snippet is left out of this version for brevity, but it contains the same functionality described in figures 1 and 2.
This class is a simple abstraction of how the model is incorporated into the compiler flow. The class is instantiated and a reference to the ML Interface (ACPOMLInterface class) is provided to enable the class to communicate with the ML framework. This framework can be active in the background (as is the case in our implementation) or be created on an as-needed basis. In either case, the reference to the interface is provided to establish communication. The model is then configured to return a specific type of result as a part of the ACPOAdvice class. Then the features to be used as input for inference are specified as input to addFeature method. Finally, to perform inference the caller would invoke the getAdvice() method and act on the result provided.
The getAdvice() method takes several simple steps, most importantly, does prepare the input data by first invoking prepareModelInput() to ensure that model inputs are provided to the model correctly and then invoking the runModel() method. The runModel method establishes communication with the ML framework, loads the model if it is not yet loaded, sends input features in, and then requests that the inference be run. If at any point in time there is an error, the process is aborted and the runModel() method will return false, subsequently causing ACPOAdvice to set to nullptr. This will be a crucial check to verify that ML inference is completed and should be checked by the caller. It is worth noting that the prepareModelInput() and runModel() methods are virtual. The reason for this is to enable derivation of this class to represent a new custom model and allow the developer to alter the flow in a way consistent with how they envision the use of the model and how they wish to handle errors.
3.5. ML APIs & Framework
An important part of the ACPOModel class is the reference to the ACPOMLInterface class. This class is the basis for the ML API in our framework and separates the actual ML model details by abstracting the underlying ML framework away from ACPOModel class, and the compiler. The code snippet is left out of this version for brevity, but it contains the same functionality described in figure 2.
The ACPOMLInterface class is virtual and it is meant to be implemented depending on how your particular ML framework needs are accessed. If it is a compiled library linked with your version of LLVM or a separate process (as was the case for this paper), this class abstracts the exchange of data between the compiler and the ML framework so that it is fundamentally irrelevant to how the ML framework actually works. To use this class, the ACPOModel invokes routines such as loadModel, freeModel, setCustomFeatures, and others to set up the model on the side of the ML framework. Then when the input data is ready, runModel is invoked to perform inference. The result is then retrieved by the ACPOModel via the getModelResult family of methods.
In our implementation, we created a derived class that communicates with an ML framework using named pipes, i.e. os.mkfifo(). This approach allows us to separate the compiler from the ML framework completely, allowing the model implementation to be on the Python side and for it to be changed without modifying the compiler on a regular basis. While at the moment the Python side of the ACPO framework is very lightweight, its separation from the compiler permits easy build-up of tools and utilities to support more models and model types without compiler changes. Later in the experimental evaluation section, we showcase the overhead of different component of ACPO in Table 1. Note that users can leverage any ML framework of their choice as long as the input/outputs match what the interface expects and in this work to showcase the versatility of ACPO, we used two different ML frameworks, i.e., PyTorch -- for model design and training and, TensorFlow -- for model deployment.
4. ACPO Models
In this section, we provide two use cases of leveraging ACPO framework on optimizing compiler passes. As we mentioned earlier, state-of-the-art approaches showcase the benefits of adapting ML models for coarse-grain and fine-grain tuning of a number of compiler optimization passes We speculate, ACPO can benefit any mid/back-end optimization/transformation pass with an internal decision process mechanism and a set of inputs/outputs. Historically, compiler developers have incrementally improved such optimization passes by means of algorithmic and heuristic approaches (Aho et al., 1986; Hall et al., 2009) and we believe leveraging ML to enhance the internal decision-making process of optimization passes merits an in-depth investigation and lays the groundwork for many future works. To this end, we choose Loop Unroll and Function Inlining as two of the most important compiler optimizations used in many state-of-the-art compilers, including LLVM.
4.1. ACPO Loop Unrolling Model
Identifying the optimal unrolling value by which a loop needs to be unrolled is known to be a hard problem with a non-polynomial complexity (Sarkar, 2000; Ashouri et al., 2018). There have been multiple works trying to automatically construct heuristics to find the optimal unrolling count (Sarkar, 2000; Lokuciejewski and Marwedel, 2009; Murthy et al., 2010; Zacharopoulos et al., 2018). A number of factors need to be examined by the decision maker to ensure the decrease in the number of control instructions, i.e., branch misses, can outweigh the increase in the number of overall instructions and eventually, the code size bloat, i.e., higher cache misses.
A new ACPOModel class, namely ACPOLUModel can be instantiated as shown under Listing 1. The textttgetAdviceML() method first retrieves a reference to the ML interface which is used to communicate with the ML framework. It then loads a model used for inference, which causes the framework to load the appropriate model as specified in the model.acpo file. The method then sets the inputs to the model and invokes the runModel method to begin inference. The inference process takes place on the ML framework, and the runModel call is a blocking call that waits until the inference completes. When it does, a successful inference will return desired results, which can be retrieved using getModelResultI methods. However, if for any reason inference fails, the result of runModel will be false, indicating that the process failed.
In the code above we instantiate an ACPOLUModel instance, which in itself creates a persistent ML interface instance to a Python-based ML Framework. The persistent interface is simply a parallel process that is launched when inference is required. However, unlike invoking blocking calls, the ML interface remains live as a separate process with which the compiler can communicate. In the case of a subsequent loop unrolling event, the process is not created anew. Instead, a reference to an existing interface is returned and the loop unrolling pass communicates with the same ML framework instance as in the previous inference step. This permits us to keep models live in memory through the compilation process. This is particularly useful in cases such as unrolling, which is often used more than once in a compiler pipeline. This way we do not have to pay the compile-time cost of restarting and closing the ML framework, or loading ML models into memory many times.
Once the ACPOLUModel is instantiated, we invoke setCustomFeatures() to specify features as inputs to the model. This fulfills the requirements of the ACPOLUModel class and ensures all inputs are provided as needed before an inference is called. The getAdvice method is then called, which in turn invokes getAdviceML which facilitates the process of inference on behalf of the compiler. The result is then decomposed into unroll-type and unroll-factor fields, which are then obeyed by the unrolling pass and applied as directed. Note that in this work we only replace the profitability checks with ML models. Decisions regarding transformation legality are left with LLVM’s legality checks in a given pass, and if those checks fail, the transformation will not proceed, nor will the profitability model be invoked.
4.2. ACPO Function Inlining
In this section, we show how to leverage ACPO to instantiate ACPOModel class for function inlining. Function inlining is one of the fundamental compiler optimizations used by many state-of-the-art compilers. In recent years, several attempts have been made to leverage ML on optimizing the pass for both code size reduction (Trofin et al., 2021; Theodoridis et al., 2022) and performance (Ashouri et al., 2022) and it was shown the optimal function inlining problem has an exponential increase in optimization space (Ashouri et al., 2018; Theodoridis et al., 2022). A new ACPOModel class, namely, ACPOFIModel can be instantiated as shown under Listing 3 .
We invoked the ACPOFIModel by adding getACPOAdvice() after the legality checks have been made. Also, we respected the high-priority overrides. We invoke ACPOFIModel object and provide the inlining information to the pass as shown in Listing 4. We instantiate an ACPOFIModel instance, which in itself creates a persistent ML interface instance to a Python-based ML Framework, similar to what we showed earlier with loop unroll pass. Once the ACPOFIModel is instantiated, we invoke setCustomFeatures() to specify features as inputs to the model. This fulfills the requirements of the ACPOFIModel class and ensures all inputs are provided as needed before an inference is called. The getAdvice method is then called, which in turn invokes getAdviceML which facilitates the process of inference on behalf of the compiler. The result is then decomposed into ”FI-ShouldInline” field, which is then obeyed by the function inlining pass and applied as directed. Note that in this work, similar to loop unroll pass, we only replace the profitability checks with ML models.
5. Experimental Evaluation
In this Section, we provide details on our experimental campaign to showcase the efficacy of both ACPOLUModel by means of cross-validation method on Polybench Subsequently, we deploy and showcase the results we reproduced with ACPOFIModel.
ACPO Loop Unrolling Model
As mentioned earlier, Loop Unroll pass has been optimized from scratch by means of adding the necessary code to connect the interface with the ML side. Additionally, a new ML model has been designed and deployed for this matter. The details of the design and testing are provided as follows. ACPOLUModel is a 5-layer Neural Net with ReLu as the activation layer and a final Softmax layer to perform the classification task to identify the optimal unrolling factor. i.e., UP.Count values, given a loop. We have designed and implemented our model in PyTorch (Paszke et al., 2017), but once we train it, we convert the model into a TensorFlow (TF) (Abadi et al., 2016) frozen model and leverage TF’s inference mechanism to do inference at deployment. As mentioned earlier in Section 3.5, users can use any ML infrastructure of choice when training and deploying models as long as the input and output between the ML APIs and the compiler side match. The input to the model is a set of 30 handcrafted features, derived from the LLVM IR to characterize the state of any given loop being transformed. The features are passed with std::vector of pairs to the ML side. Additionally, the framework supports adding other types of features, i.e., graph-based, tokens, etc., however, in this work, we leverage handcrafted features derived from LLVM IR. The features used are shown in Table 3 and taken from our super class of ACPOFeatures which contains a large set of handcrafted features defined in different scopes, i.e., Module, Function, Call Graph, and Loops, so that users can easily add and register them into their optimization use cases. The Pearson (Benesty et al., 2009) correlation analysis between the features and the target label, i.e. classes corresponding to unrolling count, is also depicted in Figure 3. The correlation values are within the range of [-1, +1], and specifically, we aimed at designing features with the highest positive/negative correlation with the target label. We did this by using features with at least a 2% correlation.
Data Collection
We used the autotuner to collect a total of 10000 different unrolling configurations from Polybench by tuning the unroll count of individual loops (UP.Count), where the set of UP.Count values are . The autotuner uses the list of loop unrolling count opportunities to determine optimal configurations by using different search techniques. Then, the compiler uses the configurations suggested by the autotuner and generates a new executable file. This process was repeated for 100 iterations to generate our training data. Note that each iteration, depending on the number of loops an application has, generates multiple data points, hence the total number of training data is significantly higher than the number of iterations of tuning rounds. The data generation process takes around 2 days. The execution time of a number of Polybench applications can be too short to allow for a reliable measurement and to this end, we define a larger input dataset for those having an execution time of less than a second. This includes: gemver, jacobi-1d, bicg, atax, mvt, durbin, trisolv, gesummv, and deriche. In this work, we define the Loop Unrolling speedup as the following:
where is the execution time of a module having multiple loops at LLVM O3’s configuration and Loop_runtime(X), the same module having a loop unrolling count configuration of X in the set of all Unrolling configurations generated by our autotuner. In this work, total program runtime and the are collected with Linux perf. All the runtime measurements are done single-threaded, and we use numactl tool to bind the workloads to a unique CPU core. All experiments are measured on an ARMv8.2-A (Kunpeng 920) architecture @ 2.6GHz on Linux. We run each benchmark three times and before each measurement run, we flush the system page cache to avoid any perturbance into the collection of our training data and we repeat the process for all measurements having more than 1% of variance between the runs to make sure they are robust. Additionally, we treat a single instance of iteration of data generation, normally having multiple loops and generating multiple samples, one set of runs meaning that we assign the Loop_speedup value to the whole set of rows generated as a result of the configuration. This way, we can filter out bad runs from good runs simply by pruning those causing a slowdown, or even having at least a certain amount of speedup values. This way ACPOLUModel only exposes to the set of quality samples and is trained to learn the relation between the given features and the target UP.Count.
Analysis of Compilation Time Overhead
We measure the compilation-time overhead of utilizing ACPOLUModel with a breakdown of each individual task, as shown in Listing 2, and we present the breakdown in Table 1. As mentioned earlier under Section 4.1, Loop Unroll pass is registered twice at O3’s pipeline. We have implemented ACPO in a way that we load our ML models once per each module and keep the models alive in the lifetime of a module to avoid a repeated loading process of the models per each instance of inference. To this end, the other contributing factor is the number of modules per application and to a lesser extent the number of loops inside a module. The number of modules determines how many times the ML model needs to be loaded, which is an expensive task, and the number of loops determines the number of times Loop Unroll pass will be activated and ACPOLUModel will be deployed. ACPOLUModel is persistent and loaded once per module and we are working on making it even more optimized, i.e., loaded once per each application having multiple modules. For example. On Polybench’s bicg, we have two modules, namely, polybench.c and bicg.c. The former has a single loop and the latter, has 8, thus, the overall overhead would be 0.857 seconds.
| Task | Each Module | Each Loop |
| Feature Collection | 0.000135 | 0.000135 |
| Setting Features for ML Model | 0.000029 | 0.000029 |
| Total ML Inference Time | 0.396148 | 0.00699 |
| Unrolling Factor Assignment | 0.000001 | 0.000001 |
| Total Time (s) | 0.396313 | 0.007155 |
ML Training and Cross Validation
The training is carried out using leave-one-out cross-validation for which we exclude one application from the training data and use it later as the test. This ensures no data leakage occurs. Measured Top-1 accuracy of the ACPOLUModel’s predictions exhibits a geometric mean of 80% accuracy with a standard deviation of 7% on Polybench. In certain applications, e.g., trmm, 2mm, etc., the top-1 accuracy is higher than the average, and in a few such as atax, it is lower. Our investigation reveals the reason lies in the imbalanced training set for the specific cases and it can be alleviated to some extent by employing a number of preprocessing and data augmentation techniques (Wei and Zou, 2019). We leave further investigation to our future works as they are outside the scope of this paper. Subsequent to establishing the high top-1 accuracy on a cross validation mode, we evaluate the performance of the ACPOLUModel by means of deploying it on Polybench and measuring the actual benefits of our ML model. Table 2 showcases the ACPOLUModel’s results when deployed on Polybench (Pouchet, 2012) on Aarch64. We observe that on average, ACPOLUModel achieved a speedup of 4.07% among the Polybench benchmarks with a best of 49.3% on Floyd-Warshall. Additionally, we observe that out of 24 benchmarks, ACPOLUModel can outperform LLVM’s O3 on 15, excluding the close call for syr2k application.
Further investigation on our obtained results reveals that Floyd-Warshall application has a seven-loop module and the added loops from polybench.c with two instances of loop unrolling pass under O3 makes it for an opportunity set of 23 loop-nest regions which can be tuned by different unroll counts. ACPO’s prediction provides for a full unroll of a couple of loops including a multiple runtime unrolling of hot loops in the kernel which matches the optimal result our autotuner found after an exhaustive search. O3 on the other hand, only unrolled two loops by their available trip count of 3 and didn’t unroll the remaining hot loops. and thus, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops.
5.1. ACPO Function Inlining Model
Similar to Loop Unrolling, Function inlining pass has been optimized by means of adding the necessary code to connect the interface with the ML side but instead of rebuilding an ML model from scratch, we decided to leverage an existing model to showcase the ease of use and the integration capabilities of ACPO when paired with an existing ML model. For this task, we use MLGOPerf (Ashouri et al., 2022) model.
5.1.1. Function Inlining Model
MLGOPerf (Ashouri et al., 2022) is a reinforcement learning model trained with Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm. The model receives as input, 13 function features including block frequency, callsite height, block counts, etc., and predicts whether or not a callsite is to be inlined. Readers can refer to the Table 2 of the reference paper for the exhaustive list of features used and the training methodology. The model has extended MLGO (Trofin et al., 2021)’s code size reduction to optimize performance instead. As shown in Listing 4, getACPOAdvice() initiates the call to collect necessary features and finally returns the Boolean value of the decision of whether or not to inline a callsite.
5.1.2. Function Inlining Generalization Results
Table 3 showcases the experimental results of the deployed ACPOFIModel when tested with CBench. As expected, the results exhibit similar behavior to MLGOPerf’s experimental results since both are leveraging the same model, albeit differently. For instance, we have gained around 27% on security_sha and inspecting the decisions made by ACPOFIModel, we notice that LLVM’s O3 did not inline the two hot callsites namely, sha_update() and sha_final() whereas in ACPO, those two are inlined. A geometric mean showcases around 2% increase in performance by incurring around 15% code size bloat on Cbench.
5.2. Combined Results
Finally, we deploy both ACPO models in combined mode and experimentally observe the benefits on Polybench and Cbench. An inspection for O3’s pipeline reveals that function inlining triggers before the first instance of loop unrolling and thus, an enhanced function inlining may benefit the subsequent passes, including loop unrolling (Ashouri et al., 2022).
| Benchmark | ACPO-LU | ACPO Combined | ||
| Size Bloat % | Speedup | Size Bloat % | Speedup | |
| 2mm | 1.903 | 0.984 | 1.903 | 0.984 |
| 3mm | 1.904 | 0.983 | 2.805 | 0.976 |
| adi | 1.0002 | 1.012 | 1.38 | 1.04 |
| atax | 0.999 | 0.985 | 1 | 0.987 |
| bicg | 1 | 1.02 | 1 | 1.022 |
| cholesky | 1.001 | 1.102 | 1 | 1.063 |
| correlation | 1.005 | 0.957 | 1.901 | 1.002 |
| covariance | 1.001 | 1.003 | 1.902 | 1.009 |
| deriche | 1.0018 | 1.078 | 1.902 | 1.115 |
| doitgen | 1.0001 | 1.011 | 1.901 | 1.013 |
| floyd-Warshall | 1.0004 | 1.493 | 1 | 0.974 |
| gemver | 1.0004 | 0.958 | 1.902 | 1.494 |
| gesummv | 1 | 1.143 | 1 | 0.975 |
| gramschmidt | 1 | 0.91 | 1 | 0.995 |
| lu | 1.001 | 0.978 | 1.9 | 1 |
| ludcmp | 1.0008 | 1.047 | 1 | 0.996 |
| mvt | 1 | 1.046 | 1 | 1.066 |
| nussinov | 2.803 | 1.171 | 1 | 1.013 |
| seidel-2d | 1.0005 | 1.0008 | 1 | 1.174 |
| symm | 1.0004 | 1.06 | 1.902 | 1.072 |
| syr2k | 1.0009 | 0.991 | 1.903 | 0.989 |
| syrk | 1.0005 | 1.002 | 1.903 | 1.036 |
| trisolv | 1 | 1.223 | 1 | 1.232 |
| trmm | 1.0004 | 0.946 | 1 | 0.975 |
| Geomean | 1.10193 | 1.040725 | 1.383 | 1.045 |
Sz%: Size bloat, Sp: Speedup. Multicolumns 1, 3, and 4 are wrt. O3 and 2, is ACPO-FI compared against MLGO (Trofin et al., 2021).
| Benchmark | ACPO-FI | wrt. MLGO | ACPO-LU | Combined | ||||
| Sz% | Sp | Sz% | Sp | Sz% | Sp | Sz% | Sp | |
| auto_bitcount | 1 | 0.999 | 1 | 1.005 | 1.001 | 1.003 | 1 | 1.001 |
| auto_qsort1 | 1 | 0.994 | 1 | 1.003 | 1 | 0.982 | 1.91 | 0.989 |
| auto_susan_c | 1 | 0.999 | 1.567 | 1.122 | 3.685 | 1.035 | 19.797 | 1.066 |
| auto_susan_e | 1 | 1.023 | 1.567 | 1.04 | 3.685 | 1.019 | 19.797 | 1.026 |
| auto_susan_s | 1 | 0.98 | 1.567 | 1.008 | 3.685 | 1.038 | 19.797 | 0.995 |
| bzip2d | 1.441 | 0.991 | 1.239 | 1.021 | 2.224 | 1.01 | 3.541 | 1.014 |
| bzip2e | 1.441 | 1.007 | 1.189 | 1.041 | 2.35 | 1.001 | 3.546 | 1.015 |
| cons_jpeg_c | 1.288 | 1.017 | 1.2 | 1.005 | 5.385 | 0.986 | 5.086 | 1.027 |
| cons_jpeg_d | 1.289 | 0.923 | 1.19 | 1.335 | 5.4 | 1.073 | 4.807 | 1.026 |
| cons_lame | 1.437 | 1.027 | 1.137 | 1.007 | 4.501 | 1.03 | 4.729 | 1.034 |
| cons_mad | 1.217 | 1.071 | 1.16 | 1 | 4.321 | 0.99 | 3.193 | 0.968 |
| cons_tiff2bw | 1.18 | 1.002 | 1.351 | 1.23 | 4.502 | 0.943 | 2.471 | 0.987 |
| cons_tiff2rgba | 1.18 | 1.25 | 1.356 | 1.11 | 4.681 | 1.036 | 2.657 | 1.012 |
| cons_tiffdither | 1.18 | 1.046 | 1.14 | 1.014 | 1.213 | 0.967 | 2.655 | 1.002 |
| cons_tiffmedian | 1.18 | 1.017 | 1.048 | 1.056 | 1 | 0.97 | 3.021 | 0.983 |
| network_dijkstra | 1 | 0.993 | 1 | 0.996 | 1 | 0.99 | 1 | 1.037 |
| network_patricia | 1 | 0.98 | 1.052 | 0.99 | 1 | 0.974 | 1 | 1.004 |
| office_ghostscript | 1.275 | 1 | 1.16 | 1.04 | 1 | 1 | 1.12 | 1.001 |
| office_ispell | 0.996 | 1 | 1.17 | 1.02 | 1 | 1 | 1 | 1 |
| office_rsynth | 1 | 1 | 1.12 | 1.04 | 2.024 | 1.01 | 2.024 | 1.011 |
| office_strings | 1 | 1.006 | 1 | 1 | 1 | 1 | 1 | 1 |
| secur_blowfish_d | 1 | 1.002 | 1 | 1.002 | 1 | 0.978 | 1 | 0.965 |
| secur_blowfish_e | 1 | 1.003 | 1.39 | 1.029 | 1 | 1.015 | 1 | 0.981 |
| security_pgp_d | 1.861 | 0.993 | 1.12 | 1 | 1 | 0.991 | 13.144 | 1 |
| security_pgp_e | 1.861 | 1.017 | 1.12 | 1.042 | 1 | 1.006 | 13.144 | 1.006 |
| secur_rijndael_d | 1 | 0.99 | 1.39 | 1.039 | 3.823 | 0.966 | 1 | 0.966 |
| secur_rijndael_e | 1 | 1.001 | 1.09 | 1.035 | 3.823 | 0.974 | 1 | 0.995 |
| security_sha | 0.999 | 1.27 | 1.009 | 1.132 | 1 | 1.174 | 0.999 | 1.501 |
| tele_adpcm_c | 1 | 1 | 0.996 | 1.001 | 1 | 1.008 | 1 | 1.038 |
| tele_adpcm_d | 1 | 0.997 | 1 | 1 | 1 | 1.132 | 1 | 1.138 |
| telecom_CRC32 | 1 | 1.094 | 1 | 1.066 | 1 | 1.042 | 1 | 1.112 |
| telecom_gsm | 1.81 | 1.076 | 1.05 | 1.04 | 1.815 | 1.15 | 1.81 | 1.117 |
| Geomean | 1.152 | 1.021 | 1.161 | 1.041 | 1.818 | 1.01 | 2.442 | 1.024 |
The combined results are shown in Tables 2 and 3. We observe a slight increase in performance benefit for both benchmarks, i.e., from 4% to 4.5% for Polybench and from 2% to 2.4% for Cbench, respectively. Note that we don’t deploy ACPO-FI on Polybench as it consists of small nested-loop kernels which might not directly benefit from an optimized inliner, however, the experimental result reveals that indirectly it can. The underlying reason is that by it can provide more opportunities for the subsequent passes registered under O3 (Ashouri et al., 2022). LLVM’s O3 pipeline has a complex behavior when it comes to pass activation and as mentioned earlier, several passes including Loop Unrolling are activated multiple times thus, it is a hard task to quantify the benefits of each model when deployed together and the accurate measurement of partial benefits, especially when optimizations are working at different scopes, i.e., Call Graph vs Function level.
One exceptionally good performing benchmark is security_sha, which we had already optimized by 27% (show on Table 3) using ACPO’s Inliner, now with 50% performance gain. There are a series of decisions made by ACPO models that we highlight an excerpt under Listing 5. Additionally, each corresponding line at Listing 6 shows the decisions made by LLVM’s O3. Note that, partial and runtime unrolling are both disabled by default at LLVM O3 and thus, F[sha_transform] Loop %for.body5 is not unrolled in Listing 6. To be abundantly fair, we measured the performance of O3 when these two flags were forced enabled as well and we noticed that the decision to unroll the loop was changed to 8 as opposed to ACPOLUModel’s prediction of 16 shown in Listing 5. However, evaluations showed that still ACPO was outperforming O3 by a margin of 7% on unrolling this particular benchmark. Further inspection revealed that O3’s threshold mechanism built into the current version of Loop Unroll pass was limiting the assignment of higher values of unroll count even though the trip count was available at compile time. Similarly, the decision to not inline several callsites including sha_transform is limited by the Inliner’s threshold mechanism.
6. Discussion
ACPO provides a class of features at varying LLVM scopes, therefore, new passes can utilize the already available features with minimal or no efforts. It also provides the flexibility of supporting other types of features, i.e., Graph-based, embeddings, etc. Additionally, we provide the Ahead-Of-Time (AOT) compilation facilities as an alternative method of inference to the developers. As mentioned in Section 3.5, ACPO uses Python named pipes to perform interprocess communication. Although this allows for greater flexibility, it adds a slight overhead to the compilation time (Shown in Table 1). Developers can switch between these modes when building the compiler. The flexibility of choosing between the two modes allows for a fast iterative prototyping of ML models during design and training without the need to rebuild the compiler infrastructure after every change and also, an efficient compilation overhead at deployment once the ML models are optimized and ready to be shipped.
7. Conclusion
In this paper, we presented ACPO, the first scalable and modular ML-guided compiler framework capable of optimizing performance for LLVM. We demonstrated that using ACPO framework, we are able to enable Loop Unroll and Function Inline passes to benefit from a seamless integration of ML models, i.e., ACPOLUModel and ACPOFIModel, for which we experimentally achieved on average 4% and %4.5 on Polybench, and 2.1% and 2.5% on Cbench. ACPO provides a generalized feature collection class for which compiler developers can add to their instantiated ACPOModel class of their need and provides utility classes to use other types of features as well. Future works will be focused on extending ACPO framework by enabling other features and optimization passes inside LLVM’s pipeline and providing innovative software characterization methods using large Languages Models (LLMs).

### v3

@@
-ACPO: AI-Enabled Compiler-Driven Program Optimization
-Abstract.
-The key to performance optimization of a program is to decide correctly when a certain transformation should be applied by a compiler.
-This is an ideal opportunity to apply machine-learning models to speed up the tuning process; while this realization has been around since the late 90s, only recent advancements in ML enabled a practical application of ML to compilers as an end-to-end framework.
-This paper presents ACPO: AI-Enabled Compiler-driven Program Optimization; a novel framework to provide LLVM with simple and comprehensive tools to benefit from employing ML models for different optimization passes.
-We first showcase the high-level view, class hierarchy, and functionalities of ACPO and subsequently, demonstrate a couple of use cases of ACPO by ML-enabling the Loop Unroll and Function Inlining passes and describe how ACPO can be leveraged to optimize other passes.
-Experimental results reveal that ACPO model for Loop Unroll is able to gain on average 4% compared to LLVM’s O3 optimization when deployed on Polybench.
-Furthermore, by adding the Inliner model as well, ACPO is able to provide up to 4.5% and 2.4% on Polybench and Cbench compared with LLVM’s O3 optimization, respectively.
-30(2,0.5)
-LLVM Dev 2023 Presentation: https://youtu.be/jzT-tar_X0U
-1. Introduction
-State-of-the-art compilers are built with sophisticated optimization pipelines that result in robust, reproducible (Ivie and Thain, 2018), and reliable code generation (Aho et al., 1986; Hall et al., 2009).
-The LLVM compiler framework is no exception with its multi-layer design, where passes in the mid-end can perform analyses or transformations on a code region, one after another (Lattner, 2008).
+ACPO: An AI-Enabled Compiler Framework
+Abstract
+The key to the performance optimization of a program is to decide correctly when a compiler should apply a certain transformation.
+This paper presents ACPO: An AI-Enabled Compiler Framework, a novel framework that provides LLVM with simple and comprehensive tools to benefit from employing ML models for different optimization passes.
+We first showcase the high-level view, class hierarchy, and functionalities of ACPO and subsequently, demonstrate a couple of use cases of ACPO by ML-enabling the Loop Unroll and Function Inlining passes used in LLVM’s O3.
+and finally, describe how ACPO can be leveraged to optimize other passes.
+Experimental results reveal that the ACPO model for Loop Unroll can gain on average 4%, 3%, 5.4%, and 0.2% compared to LLVM’s vanilla O3 optimization when deployed on Polybench, Coral-2, CoreMark, and Graph-500, respectively.
+Furthermore, by including both Function Inlining and Loop Unroll models, ACPO can provide a combined speedup of 4.5% on Polybench and 2.4% on Cbench when compared with LLVM’s O3, respectively.
+20(1.6,0.5)
+* The short version is accepted at ACM/IEEE CASES’24 – https://ieeexplore.ieee.org/document/10740757
+* LLVM-Dev 2023 Presentation: https://youtu.be/jzT-tar_X0U
+I Introduction & Related Works
+State-of-the-art compilers, e.g., LLVM, are built with sophisticated optimization pipelines that result in robust, reproducible [1], and reliable code generation [2].
+The LLVM compiler framework is no exception with its multi-layer design, where passes in the mid-end can perform analyses or transformations on a code region, one after another [3].
 LLVM passes operate on code regions with varying levels of granularity, including loops, functions, call graphs, and modules.
-Compiler engineers have continuously improved the quality of the optimization passes in LLVM, and have derived a set of fixed-length optimization pipelines, known as standard optimization levels, i.e., O{1,2,3,s,z}, which are designed to improve the code size or performance (to varying degree) of the programs to which they are applied.
+Compiler engineers have continuously improved the quality of the optimization passes in LLVM, and have derived a set of fixed-length optimization pipelines, known as standard optimization levels, i.e., O{1,2,3,s,z}, which are designed to improve the code size or performance (to varying degrees) of the programs to which they are applied.
 At the highest optimization level available, i.e.
-O3, more than 60 unique transformation passes exist in LLVM’s pipeline (Ashouri et al., 2017).
-Leveraging machine learning (ML) for compilers has become an extensively researched topic in the past years.
-In many instances, the authors showed experimental results that outperformed LLVM’s O3 (Agakov et al., 2006; Chen et al., 2012; Park et al., 2013; Ashouri et al., 2016, 2017; Cummins et al., 2017; Georgiou et al., 2018; Haj-Ali et al., 2020; Ashouri et al., 2022) or Deep Neural Networks (DNN) kernel schedulers (Patabandi et al., 2021; Patabandi and Hall, 2023) and lately, using Large Language Models (LLMs) (Cummins et al., 2023; Gong et al., 2024).
-Machine learning approaches normally rely on data generated by an autotuner (Ansel et al., 2014; Kalyan et al., 2019), which helps with exploring an optimization space formed by the number of parameters in the decision set.
+O3, more than 60 unique transformation passes exist in LLVM’s pipeline [4].
+Leveraging machine learning (ML) for compilers has recently become an extensively researched topic.
+In many instances, the authors showed experimental results that outperformed LLVM’s O3 [5, 4, 6, 7, 8, 9, 10, 11, 12].
+Machine learning approaches normally rely on data generated by an autotuner [13, 14], which helps explore an optimization space formed by the number of parameters in the decision set.
 Alternatively, by the growth of applications of Reinforcement Learning (RL), as opposed to generating offline training data by an autotuner, authors have explored their action space, online, and during the training process of the RL agent.
-This is done by means of feeding back a reward value to update the agent’s policy (Trofin et al., 2021; Cummins et al., 2021b; Ashouri et al., 2022; Wang et al., 2022a).
-To characterize the software being compiled for the ML model, program representation methods are employed, which generally fall into three categories: using static features (Fursin et al., 2011; Ashouri et al., 2022), using dynamic features (Cavazos et al., 2006; Hoste and Eeckhout, 2007), and graph-based characterization (Park et al., 2012; Ben-Nun et al., 2018; Alon et al., 2019; Cummins et al., 2021a).
-In recent years, there have been considerable advancements in the space, especially in the graph-based program representation techniques (Guo et al., 2020; Cummins et al., 2021a) and token-based identifiers (Wang et al., 2021), which are outside the scope of this work.
-The vision for ACPO is to enable the use of Machine Learning in a compiler, in a way that is modular, extendable, and easy to use.
-ACPO aims to allow for easy collaboration between compiler developers and ML architects by use of clear and simple APIs that enable the collection of relevant data for training purposes, providing inputs to an ML model during inference, and separating the compiler from the model and ML framework to enable model development and testing within the compiler context with minimal intrusion into the compiler code base.
-This enables the compiler engineers to optimize passes with an easy-to-use set of tools and limited ML knowledge.
-At the same time, once an ML model is integrated into the compiler, it can be easily optimized and redesigned with no changes to the compiler code base, effectively enabling ML architects to contribute to compiler improvement.
-This is possible due to ACPO’s separation of the ML framework and the compiler, making it possible for the compiler to work just as easily with any ML model as it is to work with any ML framework, e.g., TensorFlow (Abadi et al., 2016), PyTorch (Paszke et al., 2017), etc, and even other representation mechanism (Fursin et al., 2008; Hoste and Eeckhout, 2007; Cummins et al., 2021a).
-To this end, ACPO proposes the following contributions:
-As a framework, it provides a comprehensive set of tools, libraries, and algorithmic methods by enabling compiler engineers with a straightforward and user-friendly interface to instantiate ACPOModel classes to replace LLVM’s existing hard-coded heuristics.
+This is done by feeding back a reward value to update the agent’s policy [15, 16, 8, 17].
+To characterize the software being compiled for the ML model, program representation methods are employed, which generally fall into three categories: using static features [18, 8], using dynamic features [19, 20], and graph-based characterization [21, 22, 23, 24].
+In recent years, there have been considerable advancements in the space, especially in the graph-based program representation techniques [25, 24] and token-based identifiers [26], which are outside the scope of this work.
+Recently, there have been a few works providing an end-to-end approach to construct ML-based heuristics for optimizing the decision-making process of the transformation passes [6, 27, 28, 15, 16, 29, 8].
+Among these, only a subset provides an ML-guided optimization approach (MLGO) [16, 28, 15, 8] baked into a single or multiple passes to streamline the optimization process inside the pass functionality readily.
+The recent advancements in this area beg the question of how scalable, user-friendly, and comprehensive a framework should be to provide seamless integration with existing LLVM optimization passes while maintaining a clean separation of concerns between the ML and the LLVM layers.
+MLGO uses ML to advise the Inliner as to whether or not a call site should be inlined to minimize the size of the generated code.
+Later, MLGOPerf [8] extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on several benchmarks, e.g.
+SPEC CPU 2006 and Cbench.
+In this work, similar to MLGOPerf, we also focus on the hard problem of ML performance optimization, where the key challenge is the development of model architecture and the features used to capture key performance use cases.
+However, we develop a scalable, extendible, end-to-end ML-guided framework to enable easy use of ML within optimization passes in compilers such as LLVM.
+Contrary to MLGO, ACPO also provides a standalone and modular feature collection class defined in different scopes of LLVM, i.e., Module, Function, CallGraph, and Loops.
+Users can easily add selected features to their ACPO passes by the feature names or scopes.
+To summarize, ACPO proposes the following contributions:
+1.
+A framework to provide a comprehensive set of predefined program features, libraries, and algorithmic methods by enabling compiler engineers with a user-friendly interface to instantiate ACPOModel classes to replace LLVM’s existing hard-coded heuristics.
+2.
 ML APIs and the compiler are seamlessly connected, but at the same time, they are not interdependent; changing compiler versions, ML models, or ML frameworks won’t break the functionality of ACPO.
-As long as the inputs/outputs match, users can easily revise the ML side without the need to rebuild LLVM after every change.
-Showcasing the benefits of ACPO, we demonstrate two different scenarios:
-Loop Unroll Pass — Building both the interface and the ML model.
-Loop Unroll pass can be optimized from scratch by adding the necessary code to instantiate ACPOLUModel from ACPOModel and also by providing an ML model to outperform LLVM’s unrolling decisions at O3 on Polybench and a number of other benchmarks.
-Function Inlining Pass — Building the interface and leveraging an existing ML model.
-Function Inlining pass can be optimized by leveraging an existing ML model, i.e., MLGOPerf (Ashouri et al., 2022), using ACPO framework.
-We do that by creating an ACPOModel instance, called ACPOFIModel, and ensuring its inputs and outputs match the existing model so it can be simply invoked at the appropriate point in the compilation process.
-We demonstrate the ease of integration and versatility of ACPO framework and that how it can be used to connect existing models to the new frameworks.
-Experimental results on Cbench reveal that the performance of ACPOFIModel is comparable with the reported results in the original work while benefiting from a multi-layer design of ACPO on the separation between the ML APIs and the compiler.
+As long as the inputs/outputs match, users can easily revise the ML side without rebuilding LLVM after every change.
+3.
+Showcasing the benefits of ACPO, we demonstrate two different scenarios with LLVM:
+a) Loop Unroll Pass — Building both the interface and the ML model. b) Function Inlining Pass — Building the interface and leveraging an existing ML model [8].
 The rest of the paper is organized as follows.
-Section 2 discusses the state of the art.
-Section 3 gives details on our proposed framework.
-Under Sections 4.1 and 4.2, we showcase how Loop Unroll and Function Inlining passes can benefit by instantiating ACPOLUModel.
-Section 5 showcases the experimental results including the individual and the combined results of deploying both models and Finally, in Section 6 we discuss challenges and propose directions for future work.
-2. Related Work
-A recent survey on the compiler autotuning and optimization (Ashouri et al., 2018) reveals that there have been a vast number of approaches to employ ML techniques for automatically constructing optimization heuristics.
-These approaches mainly focused on two major problems of compiler optimization: (1) the optimization selection (Bodin et al., 1998; Cooper et al., 2002; Agakov et al., 2006; Fursin and Temam, 2009; Park et al., 2013; Ashouri et al., 2016) and (2) the phase ordering problem (Kulkarni and Cavazos, 2012; Ashouri et al., 2017; Nobre et al., 2018; Huang et al., 2019; Mammadli et al., 2020; Cummins et al., 2021b; Wang et al., 2022a).
-Further classification of the two aforementioned problems and by the type of the approach, provides insights for categorizing the works into three groups: (1) autotuning-based approaches (Fursin et al., 2011; Chen et al., 2012; Ashouri et al., 2013; Ansel et al., 2014; Kalyan et al., 2019; Cummins et al., 2021b; Wang et al., 2022b; Park et al., 2022), (2) end-to-end frameworks of optimizations (Fursin et al., 2008; Cummins et al., 2017; Haj-Ali et al., 2020; Cummins et al., 2021b; Rotem and Cummins, 2021; Das et al., 2020; Mannarswamy and Das, 2022; Ashouri et al., 2022; Wang et al., 2022b), and (3) ML-guided approaches (Rotem and Cummins, 2021; Trofin et al., 2021; Ashouri et al., 2022).
-Autotuning-based approaches leverage an iterative methodology to wrap around the compiler and to speed up the search in finding better solutions to optimizing a single or multiple passes, often constructing a new sequence of optimization to outperform standard optimization levels by means of phase-ordering or optimization selection (Agakov et al., 2006; Ashouri et al., 2016; Cummins et al., 2021b; Wang et al., 2022a).
-This can be done in a coarse (Kulkarni and Cavazos, 2012; Ashouri et al., 2016, 2017) or fine-grained manner (Kalyan et al., 2019; Phothilimthana et al., 2021; Cummins et al., 2021b).
-End-to-end approaches often provide a framework that can be paired or built with the compiler and to offer boosted search or ordering of optimizations (Cummins et al., 2017; Haj-Ali et al., 2020; Mannarswamy and Das, 2022).
+Section II discusses the state of the art.
+Section III gives details on our proposed framework and how the different layers of abstraction communicate with each other.
+Under Sections IV-A and IV-B, we showcase how Loop Unroll and Function Inlining passes can benefit by instantiating ACPOLUModel.
+Section V showcases the experimental results including the individual and the combined results of deploying both models and Finally, in Section VI we discuss challenges and propose directions for future work.
+II Related Work
+A survey on the compiler autotuning and optimization [30] reveals that there have been a vast number of approaches to employ ML techniques for automatically constructing optimization heuristics.
+These approaches mainly focused on two major problems of compiler optimization: (1) the optimization selection [31, 32, 33, 34, 35, 5] and (2) the phase ordering problem [36, 4, 37, 38, 39, 16, 17].
+Further classification of the two aforementioned problems and by the type of the approach provides insights for categorizing the works into three groups: (1) autotuning-based approaches [18, 40, 41, 13, 14, 16, 42, 43], (2) end-to-end frameworks of optimizations [44, 6, 7, 16, 28, 27, 29, 8, 42], and (3) ML-guided approaches [28, 15, 8].
+Autotuning-based approaches leverage an iterative methodology to wrap around the compiler and to speed up the search in finding better solutions to optimizing a single or multiple passes, often constructing a new sequence of optimization to outperform standard optimization levels through phase-ordering or optimization selection [33, 5, 16, 17].
+This can be done in a coarse [36, 5, 4] or fine-grained manner [14, 45, 16].
+End-to-end approaches often provide a framework that can be paired or built with the compiler to offer boosted search or ordering of optimizations [6, 7, 29].
 ML-guided approaches build the decision-making capabilities of ML models into single or multiple passes that can act as a standalone ML-enabled pass.
@@
 Cummins et al.
-(Cummins et al., 2021b) propose CompilerGym, a fine-grained RL environment to speed up the search for tuning a single or multiple compiler passes defined as action space.
-The authors leverage OpenAI’s Gym (Brockman et al., 2016) library to provide a flexible environment to tackle the selection and the phase-ordering of optimization.
-The results showcase the benefits of using their RL agent to further reduce the code size of Cbench benchmarks.
-Contrary to this work, we are proposing an ML-guided approach to provide a framework to optimize passes.
-Additionally, our objective is to improve the performance as opposed to reducing the code size of a binary.
+[16] propose CompilerGym, a fine-grained RL environment to speed up the search for tuning a single or multiple compiler passes defined as action space.
+The authors leverage OpenAI’s Gym [46] library to provide a flexible environment to tackle the selection and the phase-ordering of optimization.
+The results showcase the benefits of using their RL agent to reduce the code size of Cbench benchmarks further.
+Contrary to this work, we propose an ML-guided approach to provide a framework for optimizing passes to be used standalone or inside LLVM’s O3 pipeline.
+Additionally, our objective is to improve the performance instead of reducing a binary’s code size.
+Rotem and Cummins [28] propose a Decision Trees (DS)-based framework to infer branch probabilities for use in providing profile-guided optimizations (PGO).
+The authors use hand-crafted features for the characterization of code regions and how they are able to reap the benefits of PGO without actually running the program with instrumentation to collect profile data and recompiling.
+This work is relevant to ours since the approach builds into LLVM and once trained, it can provide PGO with certain profiling information without actually running PGO.
 Trofin et al.
-(Trofin et al., 2021) propose MLGO that uses ML to advise the Inliner as to whether or not a call site should be inlined, in order to minimize the size of the generated code.
+[15] propose MLGO that uses ML to advise the Inliner as to whether or not a call site should be inlined, in order to minimize the size of the generated code.
 MLGO framework compiles the ML model into C++ libraries using Ahead-of-Time (AOT) compilation which becomes part of the compiler at build time.
 Additionally, the authors have extended the code base to provide optimization for the register allocation problem.
-Later, MLGOPerf (Ashouri et al., 2022), extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on a number of benchmarks, e.g.
+Later, MLGOPerf [8], extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on several benchmarks, e.g.
 SPEC CPU 2006 and Cbench.
-In this work, we focus on the hard problem of ML performance optimization, where the key challenge is the development of model architecture and the features used to effectively capture key performance use cases and to facilitate the generation of robust ML models.
-In the process, we develop a scalable, extendible, end-to-end ML-guided framework to enable easy use of ML within optimization passes in compilers such as LLVM.
-We demonstrate the effectiveness and the generality of our framework by applying ML-enabling Loop Unroll and Function Inlining passes.
-We show how ML can easily and effectively be incorporated into a compiler while maintaining separation from the underlying ML frameworks.
-This framework is easily applicable to other passes and tasks within the compiler domain.
-3. Proposed Methodology
-Figure 1 presents the high-level design of our infrastructure.
-There are multiple components involved with the ACPO architecture: Autotuner, compiler, ACPOModel abstraction, low/high-level ML APIs, and ML Framework.
+This work is relevant to us as introduced and upstream the necessary libraries and foundation to train and deploy an ML model to reduce code size reduction.
+As mentioned earlier, contrary to MLGO, ACPO focuses on the hard problem of ML performance optimization.
+We show that ACPO is easily extendible and scalable to any optimization passes and we also provide a modular feature collection class, defined in different scopes of LLVM, Module, Function, CallGraph, and Loops, so users can leverage the ready-to-use set of features in their passes of interest.
+III Proposed Methodology
+Figure 2 presents the high-level flow of the design of our infrastructure.
+There are multiple components involved with the ACPO architecture: Autotuner, compiler, ACPOModel abstraction, ML APIs, and ML Framework.
 For the most part, the autotuner is a wrapper around the compiler and its hooks in the compiler are used to generate model training data.
-During the inference flow, i.e., when ML models are used for decision making within the compiler, the autotuner is not used and we can think of the compiler as being at the top of the stack.
+During the inference flow, i.e., when ML models are used for decision-making within the compiler, the autotuner is not used, and we can think of the compiler as being at the top of the stack.
 The ACPO inference flow begins when an ML-enabled pass is invoked and creates an instance of ACPOModel.
-The instance of ACPOModel aggregates appropriate features from the code region being analyzed and specifies the kind of output required from the ML model.
-The ACPOModel object then transfers the features, along with information about the model being invoked, to our ML framework via a set of remote procedure call APIs.
+The instance of ACPOModel aggregates appropriate features from the analyzed code region and specifies the kind of output required from the ML model.
+The ACPOModel object then transfers the features and information about the model being invoked to our ML framework via a set of remote procedure call APIs.
 The APIs provide the ability to send requests, including loading an appropriate model into the ML framework, executing an inference call, and returning the result of the inference back to the compiler.
 The compiler then applies transformations to the input program as prescribed in the inference output it receives.
-In the following sections, we describe in more detail the training and inference flows, and then elaborate on each component of the ACPO framework to show how they fit with the training and inference flows.
+In the following, we describe in more detail the training and inference flows, and then elaborate on each component of the ACPO framework to show how they fit with the training and inference flows.
 Autotuner
+An autotuner is a software tool that derives decisions for an optimization scenario (minimization or maximization), by intelligently exploring the design space of compiler optimization parameters [47, 13, 30, 48].
 The autotuner parameters can be either coarse- or fine-grained.
@@
 instructions, loops, functions, etc., independently of each other.
-In this work, we perform fine-grained autotuning with an autotuner (Kalyan et al., 2019), which is built on top of OpenTuner (Ansel et al., 2014).
+In this work, we perform fine-grained autotuning with an autotuner [14], which is built on top of OpenTuner [13].
 This allows us to generate meaningful training data and target labels that capture the effect of optimization parameters on code region types of interest.
-3.1. ACPO Training
+III-A ACPO Training
 A key aspect of creating an ML-driven compiler is the training of models on sample data to ensure that the compiler makes decisions consistently with good results contained in the training data.
@@
 To generate data for training purposes we rely on the autotuner.
-To do this efficiently, we enabled what we refer to as fine-grained autotuning to control each decision of a given transformation pass independently of one another.
-For example, in the case of loop unrolling, we enabled the autotuner to search for a good unroll factor for each loop, instead of applying an unroll threshold that would be specified on the compiler command line and hence applied globally on all loops.
-With the help of fine-grained autotuning, we are able to specify which pass we wish to tune and apply this kind of tuning to a set of benchmarks.
-At the compilation stage, with the instructions from the autotuner, the compiler proceeds to optimize an application and eventually reaches the ML-enabled pass for which the autotuner is currently providing decisions.Note that ACPO supports leveraging different scopes of features as long as it is supported by the compiler and it is up to the developer to decide which ACPO pass might benefit from which scope during the process of design and implementation.
-The features are passed with std::vector of pairs to the ML side.
-This process can also be scaled to an arbitrary number of features and feature types.
-The transformation decision as prescribed by the autotuner is applied to the code region (e.g.
-the autotuner can set the unroll factor of a loop to a specific value), and then the compilation continues normally until it reaches another code region that requires a decision from the ML model, at which time the process is repeated.
-Once the compilation is completed, the autotuner runs the program and collects a measurement for performance feedback (the simplest metric being the execution time of the program).
+The autotuning process allows us to override compiler heuristics with an iterative search for the optimal optimization decisions for a given application.
+By performing the search on our benchmark programs and recording each step of the search, we can obtain information about the impact of each decision on program performance, and thus collect sufficient information to be able to train the eventual model with a reasonable chance of success.
 We use the O3 compilation as our baseline and measure impact in terms of speedup relative to the O3 compilation.
-3.2. ACPO Inference
+This process is repeated many times to gather as much data per benchmark as possible to achieve beneficial results on most benchmarks.
+The resulting trial data is then processed and used as input to the model training.
+III-B ACPO Inference
 The compilation process involving ML inference is depicted in Figure 2b.
@@
 The first step in the inference process is the instantiation of the ML model, via the ACPOModel abstraction.
-This step sets up the interface to the underlying ML framework, such as TensorFlow (Abadi et al., 2016).
+This step sets up the interface to the underlying ML framework, such as TensorFlow [49].
 Upon successful framework initialization, ACPOModel collects features required for inference and sends them to the ML framework.
@@
 the predicted optimal unroll factor for a given loop.
-To accomplish the tasks of training and inference we used several key components, which we describe in greater detail in the following subsections.
-3.3. Compiler
-The compiler of choice is LLVM (Lattner, 2008) version 12, which at a high level comprises the C/C++ frontend Clang, the mid-end optimizer, and a target-specific backend to generate executable code.
-The compiler first translates a program into LLVM IR, and then proceeds to optimize the IR via a sequence of passes.
-The passes are applied in sequence and at various scopes (module, function, loop, basic block).
-Once a pass has finished transforming a code region, appropriate internal data structures, including IR, are updated and the pass either proceeds to the next code region or terminates to enable the next pass to proceed.
-Once the compilation is done, an executable is produced that can be run and measured.
-We generally refer to the sequence of passes as a pipeline, and we focus on keeping the pipeline unchanged.
-This is important because empirical evidence shows that the order of passes and composition of the compiler pipeline affect program performance (Ashouri et al., 2017).
-In this work we focus on improving decision making within a pass using ML, so changes to the pipeline ordering and optimization selection of passes are orthogonal and outside the scope of the work.
-3.4. ACPOModel
-The ACPOModel class is responsible for storing model-specific information and leveraging ML APIs to performing inference.
-The header file code snippet is left out of this version for brevity, but it contains the same functionality described in figures 1 and 2.
+III-C ACPOModel
+The ACPOModel class is responsible for storing model-specific information and leveraging ML APIs to perform inference.
+The header file code snippet is left out of this version for brevity, but it contains the same functionality described in Figure 2.
 This class is a simple abstraction of how the model is incorporated into the compiler flow.
-The class is instantiated and a reference to the ML Interface (ACPOMLInterface class) is provided to enable the class to communicate with the ML framework.
-This framework can be active in the background (as is the case in our implementation) or be created on an as-needed basis.
-In either case, the reference to the interface is provided to establish communication.
+The class is instantiated, and a reference to the ML Interface (ACPOMLInterface class) is provided to enable the class to communicate with the ML framework.
 The model is then configured to return a specific type of result as a part of the ACPOAdvice class.
-Then the features to be used as input for inference are specified as input to addFeature method.
-Finally, to perform inference the caller would invoke the getAdvice() method and act on the result provided.
-The getAdvice() method takes several simple steps, most importantly, does prepare the input data by first invoking prepareModelInput() to ensure that model inputs are provided to the model correctly and then invoking the runModel() method.
-The runModel method establishes communication with the ML framework, loads the model if it is not yet loaded, sends input features in, and then requests that the inference be run.
-If at any point in time there is an error, the process is aborted and the runModel() method will return false, subsequently causing ACPOAdvice to set to nullptr.
-This will be a crucial check to verify that ML inference is completed and should be checked by the caller.
-It is worth noting that the prepareModelInput() and runModel() methods are virtual.
-The reason for this is to enable derivation of this class to represent a new custom model and allow the developer to alter the flow in a way consistent with how they envision the use of the model and how they wish to handle errors.
-3.5. ML APIs & Framework
+Then, the features to be used as input for inference are specified as input to addFeature method.
+Finally, to perform inference, the caller would invoke the getAdvice() method and act on the result provided.
+It is worth noting that the prepareModelInput() and runModel() methods are virtual to enable derivation of this class to represent a new custom models.
+III-D ML APIs & Framework
+The snippet for the ACPOMLInterface header is shown under the Appendix section in Listing 6.
 An important part of the ACPOModel class is the reference to the ACPOMLInterface class.
-This class is the basis for the ML API in our framework and separates the actual ML model details by abstracting the underlying ML framework away from ACPOModel class, and the compiler.
+This class is the basis for the ML API in our framework and separates the actual ML model details by abstracting the underlying ML framework away from ACPOModel class and the compiler.
 The code snippet is left out of this version for brevity, but it contains the same functionality described in figure 2.
-The ACPOMLInterface class is virtual and it is meant to be implemented depending on how your particular ML framework needs are accessed.
-If it is a compiled library linked with your version of LLVM or a separate process (as was the case for this paper), this class abstracts the exchange of data between the compiler and the ML framework so that it is fundamentally irrelevant to how the ML framework actually works.
-To use this class, the ACPOModel invokes routines such as loadModel, freeModel, setCustomFeatures, and others to set up the model on the side of the ML framework.
-Then when the input data is ready, runModel is invoked to perform inference.
-The result is then retrieved by the ACPOModel via the getModelResult family of methods.
+Communication
 In our implementation, we created a derived class that communicates with an ML framework using named pipes, i.e.
 os.mkfifo().
-This approach allows us to separate the compiler from the ML framework completely, allowing the model implementation to be on the Python side and for it to be changed without modifying the compiler on a regular basis.
-While at the moment the Python side of the ACPO framework is very lightweight, its separation from the compiler permits easy build-up of tools and utilities to support more models and model types without compiler changes.
-Later in the experimental evaluation section, we showcase the overhead of different component of ACPO in Table 1.
-Note that users can leverage any ML framework of their choice as long as the input/outputs match what the interface expects and in this work to showcase the versatility of ACPO, we used two different ML frameworks, i.e., PyTorch -- for model design and training and, TensorFlow -- for model deployment.
-4. ACPO Models
+This approach allows us to completely separate the compiler from the ML framework, allowing the model implementation to be on the Python side and for it to be changed without modifying the compiler on a regular basis.
+While at the moment, the Python side of the ACPO framework is very lightweight, its separation from the compiler permits easy build-up of tools and utilities to support more models and model types without compiler changes.
+Feature Collection Class
+We have implemented a ACPOCollectFeature class having around 130 handcrafted features in total at different scopes of LLVM’s IR, from which 80 are Function level, 35s are Loops, and around 15 are Module features.
+Users can easily select features individually or by their scope when creating ACPOModel classes.
+IV ACPO Models
 In this section, we provide two use cases of leveraging ACPO framework on optimizing compiler passes.
 As we mentioned earlier, state-of-the-art approaches showcase the benefits of adapting ML models for coarse-grain and fine-grain tuning of a number of compiler optimization passes We speculate, ACPO can benefit any mid/back-end optimization/transformation pass with an internal decision process mechanism and a set of inputs/outputs.
-Historically, compiler developers have incrementally improved such optimization passes by means of algorithmic and heuristic approaches (Aho et al., 1986; Hall et al., 2009) and we believe leveraging ML to enhance the internal decision-making process of optimization passes merits an in-depth investigation and lays the groundwork for many future works.
+Historically, compiler developers have incrementally improved such optimization passes by means of algorithmic and heuristic approaches [50, 2] and we believe leveraging ML to enhance the internal decision-making process of optimization passes merits an in-depth investigation and lays the groundwork for many future works.
 To this end, we choose Loop Unroll and Function Inlining as two of the most important compiler optimizations used in many state-of-the-art compilers, including LLVM.
-4.1. ACPO Loop Unrolling Model
-Identifying the optimal unrolling value by which a loop needs to be unrolled is known to be a hard problem with a non-polynomial complexity (Sarkar, 2000; Ashouri et al., 2018).
-There have been multiple works trying to automatically construct heuristics to find the optimal unrolling count (Sarkar, 2000; Lokuciejewski and Marwedel, 2009; Murthy et al., 2010; Zacharopoulos et al., 2018).
-A number of factors need to be examined by the decision maker to ensure the decrease in the number of control instructions, i.e., branch misses, can outweigh the increase in the number of overall instructions and eventually, the code size bloat, i.e., higher cache misses.
-A new ACPOModel class, namely ACPOLUModel can be instantiated as shown under Listing 1.
-The textttgetAdviceML() method first retrieves a reference to the ML interface which is used to communicate with the ML framework.
-It then loads a model used for inference, which causes the framework to load the appropriate model as specified in the model.acpo file.
-The method then sets the inputs to the model and invokes the runModel method to begin inference.
-The inference process takes place on the ML framework, and the runModel call is a blocking call that waits until the inference completes.
-When it does, a successful inference will return desired results, which can be retrieved using getModelResultI methods.
-However, if for any reason inference fails, the result of runModel will be false, indicating that the process failed.
+IV-A ACPO Loop Unrolling Model
+Identifying the optimal unrolling value by which a loop needs to be unrolled is known to be a hard problem with a non-polynomial complexity [51, 30].
+There have been multiple works trying to automatically construct heuristics to find the optimal unrolling count [51, 52, 53, 54].
+The snippet for the ACPOLUModel is shown under the Appendix section in Listing 7 A new ACPOModel class, namely ACPOLUModel can be instantiated as shown under Listing 7.
+To leverage ML in unrolling, we invoked the ACPOLUModel after the loop unrolling legality checks had been made.
+Also, we respected the high-priority overrides.
+The highest level is set by the explicit use of unroll-count by the user, followed by the usage of #pragma clang loop unroll_count(N).
+We invoke ACPOLUModel object subsequent to these two priorities as shown in Listing 2.
 In the code above we instantiate an ACPOLUModel instance, which in itself creates a persistent ML interface instance to a Python-based ML Framework.
@@
 Decisions regarding transformation legality are left with LLVM’s legality checks in a given pass, and if those checks fail, the transformation will not proceed, nor will the profitability model be invoked.
-4.2. ACPO Function Inlining
+IV-B ACPO Function Inlining
 In this section, we show how to leverage ACPO to instantiate ACPOModel class for function inlining.
 Function inlining is one of the fundamental compiler optimizations used by many state-of-the-art compilers.
-In recent years, several attempts have been made to leverage ML on optimizing the pass for both code size reduction (Trofin et al., 2021; Theodoridis et al., 2022) and performance (Ashouri et al., 2022) and it was shown the optimal function inlining problem has an exponential increase in optimization space (Ashouri et al., 2018; Theodoridis et al., 2022).
-A new ACPOModel class, namely, ACPOFIModel can be instantiated as shown under Listing 3 .
+In recent years, several attempts have been made to leverage ML on optimizing the pass for both code size reduction [15, 55] and performance [8] and it was shown the optimal function inlining problem has an exponential increase in optimization space [30, 55].
+A new ACPOModel class, namely, ACPOFIModel can be instantiated as shown under Listing 8 .
+The snippet for the ACPOLFIModel is shown under the Appendix section in Listing 8
 We invoked the ACPOFIModel by adding getACPOAdvice() after the legality checks have been made.
 Also, we respected the high-priority overrides.
-We invoke ACPOFIModel object and provide the inlining information to the pass as shown in Listing 4.
-We instantiate an ACPOFIModel instance, which in itself creates a persistent ML interface instance to a Python-based ML Framework, similar to what we showed earlier with loop unroll pass.
-Once the ACPOFIModel is instantiated, we invoke setCustomFeatures() to specify features as inputs to the model.
-This fulfills the requirements of the ACPOFIModel class and ensures all inputs are provided as needed before an inference is called.
-The getAdvice method is then called, which in turn invokes getAdviceML which facilitates the process of inference on behalf of the compiler.
-The result is then decomposed into ”FI-ShouldInline” field, which is then obeyed by the function inlining pass and applied as directed.
-Note that in this work, similar to loop unroll pass, we only replace the profitability checks with ML models.
-5. Experimental Evaluation
-In this Section, we provide details on our experimental campaign to showcase the efficacy of both ACPOLUModel by means of cross-validation method on Polybench Subsequently, we deploy and showcase the results we reproduced with ACPOFIModel.
-ACPO Loop Unrolling Model
-As mentioned earlier, Loop Unroll pass has been optimized from scratch by means of adding the necessary code to connect the interface with the ML side.
-Additionally, a new ML model has been designed and deployed for this matter.
-The details of the design and testing are provided as follows.
-ACPOLUModel is a 5-layer Neural Net with ReLu as the activation layer and a final Softmax layer to perform the classification task to identify the optimal unrolling factor.
-i.e., UP.Count values, given a loop.
-We have designed and implemented our model in PyTorch (Paszke et al., 2017), but once we train it, we convert the model into a TensorFlow (TF) (Abadi et al., 2016) frozen model and leverage TF’s inference mechanism to do inference at deployment.
-As mentioned earlier in Section 3.5, users can use any ML infrastructure of choice when training and deploying models as long as the input and output between the ML APIs and the compiler side match.
-The input to the model is a set of 30 handcrafted features, derived from the LLVM IR to characterize the state of any given loop being transformed.
+We invoke ACPOFIModel object and provide the inlining information to the pass as shown in Listing 3.
+V Experimental Evaluation
+Note that in an ML-guided (MLGO) approach such as ACPO, optimizing a pass overrides their existing profitability checks and thus, provides a unique opportunity to run the pass in the same predefined order the sequence of passes is run at LLVM’s O3.
+Thus, all evaluations here are based on comparing the performance of O3 with the vanilla or the ML-enabled passes.
+For the same reason and except for the performance comparison with MLGO work [15] under Table V, we are not evaluating ACPO with the many other aforementioned works as their method is outside the scope of an MLGO approach.
+To this end, we first evaluate our two passes individually, and finally, a combined evaluation is presented.
+V-A ACPO Loop Unrolling Model
+We design our ACPOLUModel with a 5-layer Neural Net using ReLu as the activation layer and a final Softmax layer to perform the classification task to identify the optimal unrolling factor.
+i.e., UP.Count values, given a loop fed as input.
+The input to the model is a set of 30 handcrafted features selected from the ACPOCollectFeature class at the scope of Loops.
 The features are passed with std::vector of pairs to the ML side.
-Additionally, the framework supports adding other types of features, i.e., graph-based, tokens, etc., however, in this work, we leverage handcrafted features derived from LLVM IR.
-The features used are shown in Table 3 and taken from our super class of ACPOFeatures which contains a large set of handcrafted features defined in different scopes, i.e., Module, Function, Call Graph, and Loops, so that users can easily add and register them into their optimization use cases.
-The Pearson (Benesty et al., 2009) correlation analysis between the features and the target label, i.e.
+Additionally, the framework supports adding other types of features, i.e., graph-based, tokens, etc.
+However, we leverage handcrafted features derived from LLVM IR in this work.
+The Pearson [56] correlation analysis between the features and the target label, i.e.
 classes corresponding to unrolling count, is also depicted in Figure 3.
-The correlation values are within the range of [-1, +1], and specifically, we aimed at designing features with the highest positive/negative correlation with the target label.
-We did this by using features with at least a 2% correlation.
-Data Collection
-We used the autotuner to collect a total of 10000 different unrolling configurations from Polybench by tuning the unroll count of individual loops (UP.Count), where the set of UP.Count values are .
-The autotuner uses the list of loop unrolling count opportunities to determine optimal configurations by using different search techniques.
-Then, the compiler uses the configurations suggested by the autotuner and generates a new executable file.
-This process was repeated for 100 iterations to generate our training data.
-Note that each iteration, depending on the number of loops an application has, generates multiple data points, hence the total number of training data is significantly higher than the number of iterations of tuning rounds.
-The data generation process takes around 2 days.
-The execution time of a number of Polybench applications can be too short to allow for a reliable measurement and to this end, we define a larger input dataset for those having an execution time of less than a second.
-This includes: gemver, jacobi-1d, bicg, atax, mvt, durbin, trisolv, gesummv, and deriche.
-In this work, we define the Loop Unrolling speedup as the following:
-where is the execution time of a module having multiple loops at LLVM O3’s configuration and Loop_runtime(X), the same module having a loop unrolling count configuration of X in the set of all Unrolling configurations generated by our autotuner.
-In this work, total program runtime and the are collected with Linux perf.
+| No | Static Features | No | Static Features |
+| 1 | PartialOptSizeThreshold | 16 | NumLoadInstPerLoopNest |
+| 2 | AllowRemainder | 17 | NumStoreInstPerLoopNest |
+| 3 | UnrollRemainder | 18 | TotLoopNestInstCount |
+| 4 | AllowExpensiveTripCount | 19 | AvgNumLoadInstPerLoopNest |
+| 5 | Force | 20 | AvgNumStoreInstPerLoopNest |
+| 6 | TripCount | 21 | NumLoadInstPerLoop |
+| 7 | MaxTripCount | 22 | NumStoreInstPerLoop |
+| 8 | Size | 23 | TotLoopInstCount |
+| 9 | InitialIVValueInt | 24 | AvgNumLoadInstPerLoop |
+| 10 | FinalIVValueInt | 25 | AvgNumStoreInstPerLoop |
+| 11 | StepValueInt | 26 | IsInnerMostLoop |
+| 12 | NumPartitions | 27 | IsOuterMostLoop |
+| 13 | IndVarSetSize | 28 | MaxLoopHeight |
+| 14 | AvgStoreSetSize | 29 | TotBlocksPerLoop |
+| 15 | AvgNumInsts | 30 | IsFixedTripCount |
+Data Collection & Setup
+We used an autotuner to collect a total of 100000 different unrolling configurations from Polybench by tuning the unroll count of individual loops (UP.Count), where the set of UP.Count values are .
+In this work, we define the Loop Unrolling speedup as the fraction between the base runtime and the loop runtime.
+The total program runtime and the are collected with Linux perf.
 All the runtime measurements are done single-threaded, and we use numactl tool to bind the workloads to a unique CPU core.
-All experiments are measured on an ARMv8.2-A (Kunpeng 920) architecture @ 2.6GHz on Linux.
-We run each benchmark three times and before each measurement run, we flush the system page cache to avoid any perturbance into the collection of our training data and we repeat the process for all measurements having more than 1% of variance between the runs to make sure they are robust.
-Additionally, we treat a single instance of iteration of data generation, normally having multiple loops and generating multiple samples, one set of runs meaning that we assign the Loop_speedup value to the whole set of rows generated as a result of the configuration.
-This way, we can filter out bad runs from good runs simply by pruning those causing a slowdown, or even having at least a certain amount of speedup values.
-This way ACPOLUModel only exposes to the set of quality samples and is trained to learn the relation between the given features and the target UP.Count.
+All experiments are measured on an ARMv8.2-A architecture @ 2.6GHz on Linux.
+We run each benchmark three times, and before each measurement run, we flush the system page cache to avoid any perturbance in the collection of our training data, and we repeat the process for all measurements having more than 1% of variance between the runs to make sure they are robust.
 Analysis of Compilation Time Overhead
-We measure the compilation-time overhead of utilizing ACPOLUModel with a breakdown of each individual task, as shown in Listing 2, and we present the breakdown in Table 1.
-As mentioned earlier under Section 4.1, Loop Unroll pass is registered twice at O3’s pipeline.
+We measure the compilation-time overhead of utilizing ACPOLUModel with a breakdown of each individual task, as shown in Listing 2, and we present the breakdown in Table II.
+As mentioned earlier under Section IV-A, Loop Unroll pass is registered twice at O3’s pipeline.
 We have implemented ACPO in a way that we load our ML models once per each module and keep the models alive in the lifetime of a module to avoid a repeated loading process of the models per each instance of inference.
@@
 ML Training and Cross Validation
-The training is carried out using leave-one-out cross-validation for which we exclude one application from the training data and use it later as the test.
-This ensures no data leakage occurs.
+The training is carried out using leave-one-out cross-validation, for which we exclude one application from the training data and use it later as the test, and this ensures no data leakage occurs.
 Measured Top-1 accuracy of the ACPOLUModel’s predictions exhibits a geometric mean of 80% accuracy with a standard deviation of 7% on Polybench.
 In certain applications, e.g., trmm, 2mm, etc., the top-1 accuracy is higher than the average, and in a few such as atax, it is lower.
-Our investigation reveals the reason lies in the imbalanced training set for the specific cases and it can be alleviated to some extent by employing a number of preprocessing and data augmentation techniques (Wei and Zou, 2019).
+Our investigation reveals the reason lies in the imbalanced training set for the specific cases and it can be alleviated to some extent by employing a number of preprocessing and data augmentation techniques [57].
 We leave further investigation to our future works as they are outside the scope of this paper.
-Subsequent to establishing the high top-1 accuracy on a cross validation mode, we evaluate the performance of the ACPOLUModel by means of deploying it on Polybench and measuring the actual benefits of our ML model.
-Table 2 showcases the ACPOLUModel’s results when deployed on Polybench (Pouchet, 2012) on Aarch64.
-We observe that on average, ACPOLUModel achieved a speedup of 4.07% among the Polybench benchmarks with a best of 49.3% on Floyd-Warshall.
-Additionally, we observe that out of 24 benchmarks, ACPOLUModel can outperform LLVM’s O3 on 15, excluding the close call for syr2k application.
-Further investigation on our obtained results reveals that Floyd-Warshall application has a seven-loop module and the added loops from polybench.c with two instances of loop unrolling pass under O3 makes it for an opportunity set of 23 loop-nest regions which can be tuned by different unroll counts.
-ACPO’s prediction provides for a full unroll of a couple of loops including a multiple runtime unrolling of hot loops in the kernel which matches the optimal result our autotuner found after an exhaustive search.
-O3 on the other hand, only unrolled two loops by their available trip count of 3 and didn’t unroll the remaining hot loops.
-and thus, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops.
-5.1. ACPO Function Inlining Model
-Similar to Loop Unrolling, Function inlining pass has been optimized by means of adding the necessary code to connect the interface with the ML side but instead of rebuilding an ML model from scratch, we decided to leverage an existing model to showcase the ease of use and the integration capabilities of ACPO when paired with an existing ML model.
-For this task, we use MLGOPerf (Ashouri et al., 2022) model.
-5.1.1. Function Inlining Model
-MLGOPerf (Ashouri et al., 2022) is a reinforcement learning model trained with Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm.
-The model receives as input, 13 function features including block frequency, callsite height, block counts, etc., and predicts whether or not a callsite is to be inlined.
-Readers can refer to the Table 2 of the reference paper for the exhaustive list of features used and the training methodology.
-The model has extended MLGO (Trofin et al., 2021)’s code size reduction to optimize performance instead.
-As shown in Listing 4, getACPOAdvice() initiates the call to collect necessary features and finally returns the Boolean value of the decision of whether or not to inline a callsite.
-5.1.2. Function Inlining Generalization Results
-Table 3 showcases the experimental results of the deployed ACPOFIModel when tested with CBench.
-As expected, the results exhibit similar behavior to MLGOPerf’s experimental results since both are leveraging the same model, albeit differently.
-For instance, we have gained around 27% on security_sha and inspecting the decisions made by ACPOFIModel, we notice that LLVM’s O3 did not inline the two hot callsites namely, sha_update() and sha_final() whereas in ACPO, those two are inlined.
-A geometric mean showcases around 2% increase in performance by incurring around 15% code size bloat on Cbench.
-5.2. Combined Results
-Finally, we deploy both ACPO models in combined mode and experimentally observe the benefits on Polybench and Cbench.
-An inspection for O3’s pipeline reveals that function inlining triggers before the first instance of loop unrolling and thus, an enhanced function inlining may benefit the subsequent passes, including loop unrolling (Ashouri et al., 2022).
+| Benchmark | Top-1 Accuracy | Benchmark | Top-1 Accuracy |
+| 2mm | 88.411 % | gesummv | 80 % |
+| 3mm | 84.254 % | gramschmidt | 75.466 % |
+| adi | 76.543 % | lu | 85.971 % |
+| atax | 57.5 % | ludcmp | 85.938 % |
+| bicg | 63.295 % | mvt | 75.326 % |
+| cholesky | 86.414 % | nussinov | 79.734 % |
+| correlation | 84.091 % | seidel-2d | 80.909 % |
+| covariance | 86.42 % | symm | 79.797 % |
+| deriche | 82.481 % | syr2k | 84.858 % |
+| doitgen | 75.487 % | syrk | 82.84 % |
+| floyd-warshall | 82.864 % | trisolv | 69.225 % |
+| gemver | 86.547 % | trmm | 90.333 % |
+| Geomean: 80.196% | |||
+Table VI shows the ACPOLUModel’s results when deployed in a cross-validation mode.
+We observe that, on average, ACPOLUModel achieved a speedup of 4.07% among the Polybench benchmarks with a best of 49.3% on Floyd-Warshall.
+Further investigation reveals that the Floyd-Warshall application has a seven-loop module, which makes it for a set of 23 tunable loop-nest regions.
+ACPO’s prediction provides for a full unroll of a couple of loops, including a multiple runtime unrolling of hot loops in the kernel.
+LLVM’s O3, on the other hand, only unrolled two loops by their available trip count of 3 and didn’t unroll the remaining hot loops.
+Thus, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops.
+Generalization
+To further evaluate the generalizability of the model, we have tested ACPOLUModel on a set of heavily loop-oriented benchmarks.
+Table IV demonstrates our result for which we gain a 5.4% speedup over CoreMark, an average 3% including an 18% speedup over AMG (from Coral-2), and a 1.3% speedup on Graph500’s SSSP.
+Analyzing the results, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops.
+For instance, Coral-2’s AMG runtime unrolling of a hot loop located at hypre_CSRMatrix Module led to an increase in performance by almost 15%, which O3 failed to identify as profitable to unroll.
+| Benchmarks | O3 | ACPO | SpUp | ||
+| Time | Size | Time | Size | ||
+| CoreMark | 20.783 | 74328 | 19.723 | 74328 | 1.054 |
+| CLOMP | 142.267 | 78952 | 140.835 | 78952 | 1.010 |
+| Quicksilver | 11.647 | 256072 | 11.657 | 527328 | 0.999 |
+| STRIDE cac. | 19.707 | 72112 | 18.83 | 72112 | 1.047 |
+| STRIDE str. | 44.842 | 71928 | 44.792 | 71928 | 1.001 |
+| STREAM | 18.899 | 72992 | 18.786 | 72992 | 1.006 |
+| AMG | 3.2922 | 564192 | 2.785 | 1292568 | 1.182 |
+| LAMMPS | 2.92 | 3802424 | 2.826 | 7082568 | 1.033 |
+| BFS | 7.275 | 539264 | 7.33 | 539240 | 0.992 |
+| SSSP | 15.508 | 539832 | 15.308 | 605344 | 1.013 |
+V-B ACPO Function Inlining Model
+Instead of designing a novel ML model from scratch, we decided to leverage an existing model to showcase the ease of use and integration capabilities of ACPO when paired with an existing ML model.
+For this task, we use the MLGOPerf [8] model.
+MLGOPerf is a reinforcement learning model, and the model receives, as input, 13 function features including block frequency, callsite height, block counts, etc., and predicts whether or not a callsite should be inlined.
+Table V showcases the experimental results of using ACPO framework for the deployment of ACPOFIModel, aka MLGOPerf’s, when tested with CBench.
+As expected, the results exhibit similar behavior to MLGOPerf’s experimental results since both leverage the same model.
+V-C Combined Results
+Sz%: Size bloat, Sp: Speedup. Multicolumns 1, 3, and 4 are wrt. O3 and 2, is ACPO-FI compared against MLGO [15].
+| Benchmark | ACPO-FI | wrt. MLGO | ACPO-LU | Combined | ||||
+| Sz% | Sp | Sz% | Sp | Sz% | Sp | Sz% | Sp | |
+| auto_bitcount | 1 | 0.999 | 1 | 1.005 | 1.001 | 1.003 | 1 | 1.001 |
+| auto_qsort1 | 1 | 0.994 | 1 | 1.003 | 1 | 0.982 | 1.91 | 0.989 |
+| auto_susan_c | 1 | 0.999 | 1.567 | 1.122 | 3.685 | 1.035 | 19.797 | 1.066 |
+| auto_susan_e | 1 | 1.023 | 1.567 | 1.04 | 3.685 | 1.019 | 19.797 | 1.026 |
+| auto_susan_s | 1 | 0.98 | 1.567 | 1.008 | 3.685 | 1.038 | 19.797 | 0.995 |
+| bzip2d | 1.441 | 0.991 | 1.239 | 1.021 | 2.224 | 1.01 | 3.541 | 1.014 |
+| bzip2e | 1.441 | 1.007 | 1.189 | 1.041 | 2.35 | 1.001 | 3.546 | 1.015 |
+| cons_jpeg_c | 1.288 | 1.017 | 1.2 | 1.005 | 5.385 | 0.986 | 5.086 | 1.027 |
+| cons_jpeg_d | 1.289 | 0.923 | 1.19 | 1.335 | 5.4 | 1.073 | 4.807 | 1.026 |
+| cons_lame | 1.437 | 1.027 | 1.137 | 1.007 | 4.501 | 1.03 | 4.729 | 1.034 |
+| cons_mad | 1.217 | 1.071 | 1.16 | 1 | 4.321 | 0.99 | 3.193 | 0.968 |
+| cons_tiff2bw | 1.18 | 1.002 | 1.351 | 1.23 | 4.502 | 0.943 | 2.471 | 0.987 |
+| cons_tiff2rgba | 1.18 | 1.25 | 1.356 | 1.11 | 4.681 | 1.036 | 2.657 | 1.012 |
+| cons_tiffdither | 1.18 | 1.046 | 1.14 | 1.014 | 1.213 | 0.967 | 2.655 | 1.002 |
+| cons_tiffmedian | 1.18 | 1.017 | 1.048 | 1.056 | 1 | 0.97 | 3.021 | 0.983 |
+| net_dijkstra | 1 | 0.993 | 1 | 0.996 | 1 | 0.99 | 1 | 1.037 |
+| net_patricia | 1 | 0.98 | 1.052 | 0.99 | 1 | 0.974 | 1 | 1.004 |
+| off_ghostscript | 1.275 | 1 | 1.16 | 1.04 | 1 | 1 | 1.12 | 1.001 |
+| office_ispell | 0.996 | 1 | 1.17 | 1.02 | 1 | 1 | 1 | 1 |
+| office_rsynth | 1 | 1 | 1.12 | 1.04 | 2.024 | 1.01 | 2.024 | 1.011 |
+| office_strings | 1 | 1.006 | 1 | 1 | 1 | 1 | 1 | 1 |
+| sec_blowfish_d | 1 | 1.002 | 1 | 1.002 | 1 | 0.978 | 1 | 0.965 |
+| sec_blowfish_e | 1 | 1.003 | 1.39 | 1.029 | 1 | 1.015 | 1 | 0.981 |
+| security_pgp_d | 1.861 | 0.993 | 1.12 | 1 | 1 | 0.991 | 13.144 | 1 |
+| security_pgp_e | 1.861 | 1.017 | 1.12 | 1.042 | 1 | 1.006 | 13.144 | 1.006 |
+| sec_rijndael_d | 1 | 0.99 | 1.39 | 1.039 | 3.823 | 0.966 | 1 | 0.966 |
+| sec_rijndael_e | 1 | 1.001 | 1.09 | 1.035 | 3.823 | 0.974 | 1 | 0.995 |
+| security_sha | 0.999 | 1.27 | 1.009 | 1.132 | 1 | 1.174 | 0.999 | 1.501 |
+| tele_adpcm_c | 1 | 1 | 0.996 | 1.001 | 1 | 1.008 | 1 | 1.038 |
+| tele_adpcm_d | 1 | 0.997 | 1 | 1 | 1 | 1.132 | 1 | 1.138 |
+| tele_CRC32 | 1 | 1.094 | 1 | 1.066 | 1 | 1.042 | 1 | 1.112 |
+| telecom_gsm | 1.81 | 1.076 | 1.05 | 1.04 | 1.815 | 1.15 | 1.81 | 1.117 |
+| Geomean | 1.152 | 1.021 | 1.161 | 1.041 | 1.818 | 1.01 | 2.442 | 1.024 |
 | Benchmark | ACPO-LU | ACPO Combined | ||
@@
 | Geomean | 1.10193 | 1.040725 | 1.383 | 1.045 |
-Sz%: Size bloat, Sp: Speedup. Multicolumns 1, 3, and 4 are wrt. O3 and 2, is ACPO-FI compared against MLGO (Trofin et al., 2021).
-| Benchmark | ACPO-FI | wrt. MLGO | ACPO-LU | Combined | ||||
-| Sz% | Sp | Sz% | Sp | Sz% | Sp | Sz% | Sp | |
-| auto_bitcount | 1 | 0.999 | 1 | 1.005 | 1.001 | 1.003 | 1 | 1.001 |
-| auto_qsort1 | 1 | 0.994 | 1 | 1.003 | 1 | 0.982 | 1.91 | 0.989 |
-| auto_susan_c | 1 | 0.999 | 1.567 | 1.122 | 3.685 | 1.035 | 19.797 | 1.066 |
-| auto_susan_e | 1 | 1.023 | 1.567 | 1.04 | 3.685 | 1.019 | 19.797 | 1.026 |
-| auto_susan_s | 1 | 0.98 | 1.567 | 1.008 | 3.685 | 1.038 | 19.797 | 0.995 |
-| bzip2d | 1.441 | 0.991 | 1.239 | 1.021 | 2.224 | 1.01 | 3.541 | 1.014 |
-| bzip2e | 1.441 | 1.007 | 1.189 | 1.041 | 2.35 | 1.001 | 3.546 | 1.015 |
-| cons_jpeg_c | 1.288 | 1.017 | 1.2 | 1.005 | 5.385 | 0.986 | 5.086 | 1.027 |
-| cons_jpeg_d | 1.289 | 0.923 | 1.19 | 1.335 | 5.4 | 1.073 | 4.807 | 1.026 |
-| cons_lame | 1.437 | 1.027 | 1.137 | 1.007 | 4.501 | 1.03 | 4.729 | 1.034 |
-| cons_mad | 1.217 | 1.071 | 1.16 | 1 | 4.321 | 0.99 | 3.193 | 0.968 |
-| cons_tiff2bw | 1.18 | 1.002 | 1.351 | 1.23 | 4.502 | 0.943 | 2.471 | 0.987 |
-| cons_tiff2rgba | 1.18 | 1.25 | 1.356 | 1.11 | 4.681 | 1.036 | 2.657 | 1.012 |
-| cons_tiffdither | 1.18 | 1.046 | 1.14 | 1.014 | 1.213 | 0.967 | 2.655 | 1.002 |
-| cons_tiffmedian | 1.18 | 1.017 | 1.048 | 1.056 | 1 | 0.97 | 3.021 | 0.983 |
-| network_dijkstra | 1 | 0.993 | 1 | 0.996 | 1 | 0.99 | 1 | 1.037 |
-| network_patricia | 1 | 0.98 | 1.052 | 0.99 | 1 | 0.974 | 1 | 1.004 |
-| office_ghostscript | 1.275 | 1 | 1.16 | 1.04 | 1 | 1 | 1.12 | 1.001 |
-| office_ispell | 0.996 | 1 | 1.17 | 1.02 | 1 | 1 | 1 | 1 |
-| office_rsynth | 1 | 1 | 1.12 | 1.04 | 2.024 | 1.01 | 2.024 | 1.011 |
-| office_strings | 1 | 1.006 | 1 | 1 | 1 | 1 | 1 | 1 |
-| secur_blowfish_d | 1 | 1.002 | 1 | 1.002 | 1 | 0.978 | 1 | 0.965 |
-| secur_blowfish_e | 1 | 1.003 | 1.39 | 1.029 | 1 | 1.015 | 1 | 0.981 |
-| security_pgp_d | 1.861 | 0.993 | 1.12 | 1 | 1 | 0.991 | 13.144 | 1 |
-| security_pgp_e | 1.861 | 1.017 | 1.12 | 1.042 | 1 | 1.006 | 13.144 | 1.006 |
-| secur_rijndael_d | 1 | 0.99 | 1.39 | 1.039 | 3.823 | 0.966 | 1 | 0.966 |
-| secur_rijndael_e | 1 | 1.001 | 1.09 | 1.035 | 3.823 | 0.974 | 1 | 0.995 |
-| security_sha | 0.999 | 1.27 | 1.009 | 1.132 | 1 | 1.174 | 0.999 | 1.501 |
-| tele_adpcm_c | 1 | 1 | 0.996 | 1.001 | 1 | 1.008 | 1 | 1.038 |
-| tele_adpcm_d | 1 | 0.997 | 1 | 1 | 1 | 1.132 | 1 | 1.138 |
-| telecom_CRC32 | 1 | 1.094 | 1 | 1.066 | 1 | 1.042 | 1 | 1.112 |
-| telecom_gsm | 1.81 | 1.076 | 1.05 | 1.04 | 1.815 | 1.15 | 1.81 | 1.117 |
-| Geomean | 1.152 | 1.021 | 1.161 | 1.041 | 1.818 | 1.01 | 2.442 | 1.024 |
-The combined results are shown in Tables 2 and 3.
+An inspection of O3’s pipeline reveals that function inlining triggers before the first instance of loop unrolling, and thus, an enhanced function inlining may benefit the subsequent passes, including loop unrolling.
+Note that there are a number of passes in between the two instances of our models, i.e., DCE, etc, which change the code size of an IR, and thus, the combined code size bloat does not reflect the addition of the two instances of deployments.
+The combined results are shown in Tables VI and V.
 We observe a slight increase in performance benefit for both benchmarks, i.e., from 4% to 4.5% for Polybench and from 2% to 2.4% for Cbench, respectively.
-Note that we don’t deploy ACPO-FI on Polybench as it consists of small nested-loop kernels which might not directly benefit from an optimized inliner, however, the experimental result reveals that indirectly it can.
-The underlying reason is that by it can provide more opportunities for the subsequent passes registered under O3 (Ashouri et al., 2022).
-LLVM’s O3 pipeline has a complex behavior when it comes to pass activation and as mentioned earlier, several passes including Loop Unrolling are activated multiple times thus, it is a hard task to quantify the benefits of each model when deployed together and the accurate measurement of partial benefits, especially when optimizations are working at different scopes, i.e., Call Graph vs Function level.
-One exceptionally good performing benchmark is security_sha, which we had already optimized by 27% (show on Table 3) using ACPO’s Inliner, now with 50% performance gain.
-There are a series of decisions made by ACPO models that we highlight an excerpt under Listing 5.
-Additionally, each corresponding line at Listing 6 shows the decisions made by LLVM’s O3.
-Note that, partial and runtime unrolling are both disabled by default at LLVM O3 and thus, F[sha_transform] Loop %for.body5 is not unrolled in Listing 6.
-To be abundantly fair, we measured the performance of O3 when these two flags were forced enabled as well and we noticed that the decision to unroll the loop was changed to 8 as opposed to ACPOLUModel’s prediction of 16 shown in Listing 5.
+Note that we don’t deploy ACPO-FI on Polybench as the benchmark consists of small kernels of nested loops, which might not directly benefit from an optimized inliner.
+However, the combined experimental result reveals that it can be done indirectly.
+LLVM’s O3 pipeline has a complex behaviour when it comes to pass activation, and as mentioned earlier, several passes, including Loop Unrolling, are activated multiple times; thus, it is a hard task to quantify the benefits of each model when deployed together and the accurate measurement of partial benefits, especially when optimizations are working at different scopes, i.e., Call Graph vs Function level.
+One exceptionally good performing benchmark is security_sha, which we had already optimized by 27% (shown on Table V) using ACPOFIModel, now with 50% performance gain.
+There are a series of decisions made by ACPO models that we highlight an excerpt under Listing 4.
+Additionally, each corresponding line at Listing 5 shows the decisions made by LLVM’s O3.
+Note that partial and runtime unrolling are both disabled by default at LLVM O3 and thus, F[sha_transform] Loop %for.body5 is not unrolled in Listing 5.
+To be abundantly fair, we measured the performance of O3 when these two flags were forced enabled as well and we noticed that the decision to unroll the loop was changed to 8 as opposed to ACPOLUModel’s prediction of 16 shown in Listing 4.
 However, evaluations showed that still ACPO was outperforming O3 by a margin of 7% on unrolling this particular benchmark.
 Further inspection revealed that O3’s threshold mechanism built into the current version of Loop Unroll pass was limiting the assignment of higher values of unroll count even though the trip count was available at compile time.
-Similarly, the decision to not inline several callsites including sha_transform is limited by the Inliner’s threshold mechanism.
-6. Discussion
-ACPO provides a class of features at varying LLVM scopes, therefore, new passes can utilize the already available features with minimal or no efforts.
+Similarly, the decision to not inline several callsites including sha_transform is limited by the LLVM Inliner’s threshold mechanism.
+VI Discussion
+ACPO is the first scalable framework for bringing ML into compilers.
+It provides the ability to reproduce models on different architecture very fast and the methodology lays out the steps to regenerate data, retrain a new model, and easily deploy with LLVM.
+Compiler autotuning and optimization space have benefited from ML for almost two decades.
+As we pointed out earlier, there are many recent works leveraging ML, however, the goal of ACPO is to provide a novel flexible framework for the LLVM community to leverage ML on any optimization passes of interest with minimal efforts.
+Nevertheless, there are a number of challenges and shortcomings we would like to mention here.
+Integration and dependencies
+ACPO attempts to alleviate the challenging task of supply chain management and the use of licensed tools by separating the ML framework from the compiler.
+We showed both the integration of PyTorch and TensorFlow among other moving objects, i.e., Autotuner, into our framework, however, upstreaming would remain challenging and requires community efforts to shape the next generations of ML-enabled compilers.
+ACPO provides a class of features at varying LLVM scopes, therefore, new passes can utilize the already available features with minimal or no effort.
 It also provides the flexibility of supporting other types of features, i.e., Graph-based, embeddings, etc.
 Additionally, we provide the Ahead-Of-Time (AOT) compilation facilities as an alternative method of inference to the developers.
-As mentioned in Section 3.5, ACPO uses Python named pipes to perform interprocess communication.
-Although this allows for greater flexibility, it adds a slight overhead to the compilation time (Shown in Table 1).
+As mentioned in Section III-D, ACPO uses Python named pipes to perform interprocess communication.
+Although this allows for greater flexibility, it adds a slight overhead to the compilation time (Shown in Table II).
 Developers can switch between these modes when building the compiler.
 The flexibility of choosing between the two modes allows for a fast iterative prototyping of ML models during design and training without the need to rebuild the compiler infrastructure after every change and also, an efficient compilation overhead at deployment once the ML models are optimized and ready to be shipped.
-7. Conclusion
-In this paper, we presented ACPO, the first scalable and modular ML-guided compiler framework capable of optimizing performance for LLVM.
-We demonstrated that using ACPO framework, we are able to enable Loop Unroll and Function Inline passes to benefit from a seamless integration of ML models, i.e., ACPOLUModel and ACPOFIModel, for which we experimentally achieved on average 4% and %4.5 on Polybench, and 2.1% and 2.5% on Cbench.
+Multi-objective optimization
+ACPO optimizes the performance of a code region with intelligent unrolling decisions.
+However, in a number of cases, we experimentally observe that the emitted code size is also increased.
+This is a(n) (in)direct result of Inliner and Loop Unroll which increase the number of instructions and thus, the code size.
+Although this is outside the scope of the current work, an enhanced version may also benefit from taking into account both of the objectives in its exploration strategy by identifying the Pareto optimality [58, 41, 59, 60].
+Similar strategies are employed for a number of adjacent problems; i.e, Multicore Embedded Systems [61, 62] and Compiler autotuning [58, 41, 59].
+VII Conclusion
+In this paper, we presented ACPO, a scalable and modular ML-guided compiler framework capable of optimizing performance for LLVM.
+We demonstrated that using the ACPO framework, we are able to enable Loop Unroll and Function Inline passes to benefit from a seamless integration of ML models, i.e., ACPOLUModel and ACPOFIModel, for which we experimentally achieved on average 4% and %4.5 on Polybench, and 2.1% and 2.5% on Cbench.
 ACPO provides a generalized feature collection class for which compiler developers can add to their instantiated ACPOModel class of their need and provides utility classes to use other types of features as well.
-Future works will be focused on extending ACPO framework by enabling other features and optimization passes inside LLVM’s pipeline and providing innovative software characterization methods using large Languages Models (LLMs).